### v1

# Rewarding Doubt: A Reinforcement Learning Approach to Confidence Calibration of Large Language Models

Abstract
A safe and trustworthy use of Large Language Models (LLMs) requires an accurate expression of confidence in their answers. We introduce a novel Reinforcement Learning (RL) approach for LLM calibration that fine-tunes LLMs to elicit calibrated confidence estimations in their answers to factual questions. We model the problem as a betting game where the model predicts a confidence score together with every answer, and design a reward function that penalizes both over and under-confidence. We prove that under our reward design an optimal policy would result in a perfectly calibrated confidence estimation. Our experiments demonstrate significantly improved confidence calibration and generalization to new tasks without re-training, indicating that our approach teaches a general confidence awareness. This approach enables the training of inherently calibrated LLMs.
Machine Learning, ICML
1 Introduction
In human intelligence and inter-human interaction, the ability to understand our own uncertainty and communicate our doubts to others is fundamental for effective decision-making, collaboration, and learning (Cosmides & Tooby, 1996; Xiong et al., 2024). Similarly, for Large Language Models (LLMs) to be safely used in real-world applications, especially when humans and AI systems work together, they must not only generate accurate information but also communicate their confidence in that information. While LLMs have demonstrated impressive capabilities in natural language understanding, question answering and text summarization (Touvron et al., 2023; Chiang et al., 2023; Achiam et al., 2023), LLMs still face significant limitations, such as their tendency to generate inaccurate information, often referred to as hallucinations (Hadi et al., 2023). This raises concerns about their reliability, particularly in real-world applications where trustworthiness is essential. Especially in high-stakes environments such as medical diagnosis, where LLMs are starting to become support tools for professionals (Moor et al., 2023; Pellegrini et al., 2023; Tu et al., 2024; Bani-Harouni et al., 2024), overconfident predictions including factual errors or hallucinations could have serious consequences for patient health. Also, in customer service or legal consultation (Shi et al., 2024; Sun et al., 2024), LLMs need to express uncertainty and defer complex queries to human representatives when unsure to avoid misinformed decisions. Reliable confidence estimation would enable these systems to flag uncertain outputs for human review, ensuring that crucial decisions are not made based on uncertain LLM outputs. To allow risk estimation while using LLM-generated output, model confidence should be calibrated, meaning that the expressed numerical confidence should be equal to the probability of the model’s answer being correct.
Figure 1: Illustration of our reward design: The model is rewarded for a high confidence if it is correct and punished if it is incorrect. To maximize the reward, the model needs to learn when to predict a higher or lower confidence, considering a possible higher punishment. Our reward function is designed so that the best reward is given when the confidence and the epistemic probability of being correct are the same, thus incentivizing the expression of calibrated confidences.
Many previous methods for confidence estimation lack in calibration performance as they do not train the model and instead infer the confidence from the internal state in a zero-shot setup (Huang et al., 2023b; Kuhn et al., 2023; Duan et al., 2024). Additionally, this does not give models an inherent awareness of confidence. Other trained methods in this area generally do not explicitly optimize for calibration but instead for a binary correctness target (Azaria & Mitchell, 2023; Kapoor et al., 2024). While this ensures strong performance in terms of discrimination, it overlooks the practical importance of fine-grained and well-calibrated confidence scores. Without proper calibration, users cannot reliably interpret confidence values, making it difficult to determine when to trust a model’s predictions. This is particularly problematic in real-world applications where decision thresholds vary based on context. For example, a high-stakes clinical decision requires high confidence, while a low-stakes trivia question might tolerate lower confidence. By contrast, well-calibrated models enable more flexible and trustworthy human interaction, allowing users to dynamically adjust their confidence requirements based on the stakes of a given decision.
Targeting these limitations, we propose a novel reinforcement-learning (RL) approach for teaching LLMs to express their calibrated confidence, encouraging a granular, accurate estimation of the confidence level in the training objective. For this, we model confidence estimation as a betting game: a high-confidence answer would warrant a larger bet, reflecting a strong belief in its correctness, while a lower confidence score would suggest caution. Our reward function is designed to reflect the risk-reward tradeoff in betting, as illustrated in Figure 1. It increases the reward when a correct answer is given with high confidence, simulating the higher stakes of big bets. Conversely, it penalizes incorrect answers more when they are made with high confidence, discouraging overconfidence. This ensures that both uncertainty and confidence are appropriately factored into the reward system. With our reward design, the model learns to align its predicted confidence with the accuracy of its output, encouraging granular and calibrated confidence scoring, as a calibrated confidence estimation will provably result in the highest reward during training. This not only improves the trustworthiness of LLMs in collaborative human-AI scenarios but also helps users better assess when AI tools should be trusted, double-checked, or deferred to human expertise.
Figure 2: Overview of our reinforcement learning framework: The LLM is prompted to answer a question and provide the confidence in this answer. The answer is checked for correctness by a judge function and the reward is calculated based on the correctness and the confidence. Correct answers with high confidences are rewarded highly, but also penalized heavily when incorrect.
2 Related Works
Confidence estimation and calibration have a long history in machine learning and natural language processing (Wang, 2024). With the rise of LLMs, a lot of attention has moved to extracting and improving confidence estimation from these new models. There is a wide range of methods that can roughly be divided into black-box and white-box methods (Geng et al., 2024).
Black-box methods require only the model generations for uncertainty estimation and do not depend on any knowledge of the model’s internal state. Within this group, linguistic confidence methods assess the uncertainty by simply prompting the LLM to verbalize its confidence. Xiong et al. (2024) showed that prompting techniques like chain-of-thought (CoT) (Wei et al., 2022) can, to a limited extend, help the model in the calibration of its confidence expression. Based on the assumption that uncertainty in its generation would prime the model to hallucinate information, consistency-based methods (Manakul et al., 2023; Wang et al., 2022) measure the similarity between multiple generations and show that high variance in the generation correlates with high uncertainty. Wang et al. (2022) propose using CoT to produce various reasoning pathways towards an answer. By marginalizing out the reasoning pathways, they provide the most consistent answer as the one most probably correct and offer the consistency measure as an estimation of the model’s confidence. Black-box methods are valuable for their simplicity, ease-of-use and universality, however generally lack behind white-box methods in their calibration performance.
White-box methods, on the other hand, depend on access to the internal state of the model, either for the evaluation of logits and activations or for fine-tuning of the model for calibration. Logit-based methods calibrate the confidence by assessing the probabilities or the entropy of tokens in the predicted sequence. Huang et al. (2023a) calculate the mean probability of the generated tokens with the intuition that a model that generates an answer with high probability tokens also has a high confidence in this answer. Kuhn et al. (2023) and (Duan et al., 2024) expand on this idea by looking at the entropy of the generated tokens while also taking the semantic relevance into account. Kadavath et al. (2022) showed that LLMs can self-evaluate the correctness of their own generation. They prompt the model to provide an answer followed by a judgement whether its own answer is "true" or "false". They then compare the probability of the "true" or "false" token to calculate a confidence estimation.
While all the aforementioned methods are zero-shot methods, another line of white-box methods, proposes to train external models for confidence estimation. Xiong et al. (2024) propose a method that trains a linear probe for assessing the model’s confidence based on the activations of the LLM’s hidden state. They evaluate the application of the probe to various layers and train it with correctness labels of the task. This approach does not explicitly train on the granularity of confidence estimation and, therefore, is only indirectly able to express a confidence gradient for fine-grained calibration. Building on this approach, Mielke et al. (2022) give an LLM access to the confidence extracted by the external probe and fine-tune it to express this confidence. However, the calibration performance is upper-bounded by the ability of the trained probe to assess the model’s confidence accurately.
Fine-tuning LLMs using Reinforcement Learning from human feedback (RLHF) has received increasing attention for aligning the models with human preferences, avoiding toxicity or the generation of harmful content (Ouyang et al., 2022; Lu et al., 2022). Outside of RLHF, RL has been used to fine-tune LLMs for agentic interaction in textual environments (Zhou et al., 2023; Carta et al., 2023).
Confidence estimation of LLMs using RL is a novel research direction that, to the best of our knowledge, has only been approached by Tao et al. (2024) and Xu et al. (2024). Tao et al. (2024) build on the Reinforcement Learning with Human Feedback (RLHF) paradigm and design a reward function that aligns the verbalized confidence of the model with the preference rating predicted by the RLHF reward model. However, this approach can not evaluate model confidence in factual statements and depends on human-annotated preference data. Xu et al. (2024) introduce a novel Reinforcement Learning from Knowledge Feedback (RLKF) framework, which aims to teach a model to refuse to answer questions outside their knowledge scope, reducing hallucinations. However, this does not quantify confidence when giving an answer.
3 Rewarding Doubt
We propose a novel reinforcement-learning approach, that improves an LLM’s ability to verbalize an accurate numerical confidence in a previously generated answer. The LLM functions as an agent in a simulated environment as shown in Figure 2, that poses challenging question-answering scenarios. It is prompted with task queries such as factual questions and asked to predict both an answer to the query as well as a confidence score. Based on the correctness of the answer, and the expressed confidence, we reward the model, incentivizing it to express a calibrated confidence.
Formally, let the model be provided with a textual question or request $q$, prompting an answer-confidence pair $(a,\hat{p})$ as response, where $a$ is a textual answer with binary correctness value, and $0\leq\hat{p}\leq 1$ is a numerical confidence score representing the subjective probability the model assigns to answer $a$ being correct. We train this subjective probability assessment to align with the true epistemic probability $p^{*}$, which represents the actual likelihood of correctness given the model’s internal knowledge state. If $\hat{p}$ and $p^{*}$ are aligned the model is perfectly calibrated, meaning the probability of correctness $P(j(a)=1)$ always equals the expressed confidence:
$P(j(a)=1\mid\hat{p}=x)=x,\quad\forall x\in[0,1],$
where $j(\cdot)$ is a correctness judging function that is 1 if answer $a$ is correct, and 0 otherwise.
The true epistemic probability $p^{*}$ is not directly observable, thus supervised learning of calibration is only possible by constructing an artificial ground truth to approximate $p^{*}$. Instead, we model this task as a Markov Decision Process (MDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},R)$, where the model learns to generate calibrated confidence scores through reinforcement learning. The MDP is defined by the following components:
  • State space ($\mathcal{S}$): A state $s_{t}\in\mathcal{S}$ consists of a natural language question $q$, the model’s predicted answer $a$, and the partial sequence of confidence tokens predicted so far, if any. That is, $s_{t}=(q,a,c_{1:t-1})$, where $c_{1:t-1}$ represents the previously generated confidence score tokens.
  • Action space ($\mathcal{A}$): The action space consists of selecting the next token $c_{t}$ in the confidence estimation process from the LLM vocabulary, including numerical tokens (e.g., representing percentages or probability values) and a special end-of-sequence token that finalizes the prediction.
  • Transition function ($\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$): The environment transitions deterministically based on the language model’s autoregressive token generation process. Given a state $s_{t}=(q,a,c_{1:t-1})$ and an action $c_{t}$, the next state is defined as $s_{t+1}=(q,a,c_{1:t})$. Once the end-of-sequence token is generated, the episode terminates.
  • Reward function ($R$): The reward $R(a,c,j)$ is computed based on the final confidence score sequence $c=(c_{1},\dots,c_{T})$ and the correctness of the answer $j(a)$.
To promote accurate confidence estimation, the model’s reward function must fulfill the requirement of being maximized when $\hat{p}=p^{*}$. The model should receive a higher reward when it correctly predicts an outcome with high confidence. Conversely, incorrect predictions should be penalized more severely when the model exhibits high confidence. In contrast, expressing uncertainty reduces potential penalties, leading to a higher expected reward in uncertain scenarios. This approach incentivizes the model to express high confidence only in cases where certainty is warranted while expressing doubt in ambiguous situations. By penalizing both overconfidence and underconfidence, the model is encouraged to calibrate its confidence accurately, effectively balancing the trade-off between reward maximization and penalty avoidance.
We design our reward as a logarithmic scoring function, fulfilling the described optimality conditions:
$R(a,\hat{p},j)=\begin{cases}log(\hat{p}),&\text{if $j(a)$ = 1 (correct)}\\ log(1-\hat{p}),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
In the following, we prove that $R(a,\hat{p},j)$ is maximized when $\hat{p}=p^{*}$ and thus optimizes for a perfect calibration:
Proof.
Let $f(\hat{p})$ be the expected reward for all values of $\hat{p}$ and $p^{*}$:
$f(\hat{p})=p^{*}\log(\hat{p})\;+\;(1-p^{*})\log\bigl{(}1-\hat{p}\bigr{)}.$
Taking the first derivative w.r.t. $\hat{p}$:
$f^{\prime}(\hat{p})=\frac{p^{*}}{\hat{p}}\;-\;\frac{1-p^{*}}{1-\hat{p}}\quad$
and setting
$f^{\prime}(\hat{p})=0\implies p^{*}(1-\hat{p})=\hat{p}(1-p^{*})\ \;\;\Longrightarrow\;\;\hat{p}=p^{*}$
showing the only critical point in $(0,1)$ of $f^{\prime}$ is at $\hat{p}=p^{*}$.
The second derivative:
$f^{\prime\prime}(\hat{p})=-\frac{p^{*}}{\hat{p}^{2}}\;-\;\frac{1-p^{*}}{\bigl{(}1-\hat{p}\bigr{)}^{2}}$
is strictly negative for $\hat{p}\in(0,1)$. Hence, $f(\hat{p})$ is concave and has its global maximum at $\hat{p}=p^{*}$. ∎
Since the logarithm of zero is undefined, we set the actual lower limit of the confidence $\hat{p}$ to 0.001 and the upper limit to 0.999. The normalized reward for correct and incorrect answers for each confidence is visualized in Figure 3. Our reward function rewards incorrect responses with a confidence of zero equally to correct responses with a maximum confidence. That means the model could exploit the reward function by always responding with a wrong answer and a confidence of zero. To counteract this possible scenario, we add a small constant value to the reward function if the answer is correct. This does not change the location of the maximum of the reward function and thus does not alter the optimality but discourages the model from systematically responding incorrectly.
Figure 3: The rewards for each confidence value for correct and incorrect answers. The closer the confidence is to ten or zero, respectively, the higher is the reward. At the same time, the possible punishment increases to a greater extent. The model has to learn when the trade-off between those two possibilities is worthwhile.
4 Experimental Setup
We evaluate our method in both Single-Answer and Multiple-Answer settings. We prompt the model to provide a confidence for each answer as an integer between 0 and 10, which we normalize for the reward calculation. A confidence of 0 is defined as the model being certain that the answer is incorrect, while 10 is defined as the model being certain the answer is correct. We normalize the reward function to the range of $[-1,1]$.
In the Single-Answer setting we train the model on the TriviaQA dataset (Joshi et al., 2017), which contains question-answer-evidence triplets, from which we only use the questions and answers. For generalization experiments, we evaluate our method on CommonsenseQA (Talmor et al., 2019) and MedQA (Jin et al., 2020), which are multiple-choice question datasets in the commonsense and medical domain, respectively. For the Multiple-Answer setting, we train on the QAMPARI dataset (Amouyal et al., 2023), which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought (Xiong et al., 2024), Surrogate Token (Kadavath et al., 2022), Sequence Probability (Huang et al., 2023b), Self-Consistency (Wang et al., 2022) and Trained Probe (Azaria & Mitchell, 2023). We also compare to the non-finetuned base model in a zero-shot manner, using the same prompt as our Rewarding Doubt method and refer to this setup as Verbalize. In the Multiple-Answer setting we compare only to Trained Probe and Sequence probability, as Chain-of-Tought and Self-Consistency have a high runtime in the Multiple-Answer setting and under-performed in the Single-Answer setting.
We report our results using the Expected Calibration Error (ECE) and the Area Under the Receiver Operating Characteristic Curve (AUROC) metric. The ECE measures the difference between a model’s predicted confidence and its actual accuracy, assessing how well confidences align with real outcomes. AUROC assesses a model’s ability to distinguish between classes, with a score of 0.5 indicating random guessing and 1.0 indicating perfect classification. Additionally, we visualize the calibration with calibration curves, where a well-calibrated model lies close to the 45° line and large deviations show a high miscalibration.
4.1 Response Generation
To calibrate and reward the model only on the confidences and not the answers we separate generation in two steps. In the Single-Answer-Setting the model is prompted to generate responses in the format "Answer: <answer>, Confidence: <confidence>". In the first generation step an additional stopping criterion is the "Confidence" token, ensuring the model stops before generating the confidence. This response is then used as input to generate the confidence. For multiple answers, the model is prompted to list all its answers together with a confidence score, where each answer-confidence pair has the same format as in the Single-Answer setting. The model first generates all answers and their corresponding confidences. Individual answers and confidences are parsed with regular expressions, and new queries to generate a confidence for each answer are created by appending tokens preceding the previous confidence token.
4.2 Correctness Assessment
For the multiple-choice datasets MedQA and CommonsenseQA, we evaluate correctness using the exact string matching between the model’s response and the ground truth answer.
For the TriviaQA and QAMPARI datasets, we use the F1 score of word overlap to measure the similarity between the model’s response and the ground truth candidates. The F1 score is calculated for each candidate and the maximum score is considered the final score. We consider an answer as correct if its score exceeds a threshold of 0.5.
4.3 Implementation Details
4.3.1 Training
As base model for our experiments we use Meta-Llama-3-8B-Instruct (Grattafiori et al., 2024). We employ the 4-bit quantized performance-optimized model version by Unsloth AI (Han et al., 2023). Additionally, we apply LoRA fine-tuning (Hu et al., 2022), reducing the amount of trainable parameters from 8,030,265,345 to 41,947,137. For the Single-Answer setting we train the model for two epochs with a learning rate of 1e-5. For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 32,000 steps with a batchsize of eight and a learning rate of 1e-5 with a reward scale of five. On average the model generated approximately 3.4 answers per fact. If the model fails to generate an answer in the specified format, it is penalized with an out-of-format reward of -3. For a correct answer we additionally reward the model with 0.25.
4.3.2 Baselines
For the Sequence Probability, we compute the average probability for each token in the response. In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric (Zhang et al., 2019). For the trained probe (Azaria & Mitchell, 2023), the original study introduced a custom dataset comprising short statements classified as either true or false. The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements. To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers. The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework. For the architecture of the MLP, we employed the same design as Azaria & Mitchell (2023) and train it for four epochs with a learning rate of 1e-4 until convergence. The exact prompts used for each baseline are provided Appendix A.
5 Results and Discussion
This section presents and discusses the key findings of our experiments.
Rewarding Doubt improves calibration significantly across all settings
To evaluate how well our approach improves calibration, we compare it against the zero-shot LLM (Verbalize) and existing methods in both the Single-Answer and Multiple-Answer settings. The comparisons for the Single-Answer setting are provided in Table 1, while those for the Multiple-Answer setting are shown in Table 2. Across all settings, our approach significantly improves the LLM’s confidence calibration compared to zero-shot verbalization and outperforms most baselines, which often completely fail to produce accurate confidence estimates, in both ECE and AUROC. The only exception is the TriviaQA dataset, where the Trained Probe achieves a slightly better ECE score, with both methods achieving near-perfect results. However, our approach achieves a higher AUROC score indicating that its confidences are more effective for distinguishing between correct and incorrect answers compared to the Trained Probe. Since the Trained Probe is the only other trained method, it is not surprising that it outperforms the non-trained baselines. These results further support the claim by Azaria & Mitchell (2023) that the model’s internal state holds information about the truthfulness of statements, which can serve as an indicator of uncertainty. However, the non-fine-tuned model appears to be unable to effectively associate this internal state with uncertainty. Our approach enables the model to make use of this correlation, resulting in a more accurate expression of the probability that a given answer is correct. Notably, our approach consistently outperforms all baselines in the AUROC metrics, even though our reward function only emphasizes calibration accuracy, whereas the Trained Probe is optimized for AUROC. As illustrated in Figure 4, the Rewarding Doubt method leads to much improved calibration curves compared to the base model for both TriviaQA and QAMPARI. The main outlier on QAMPARI at the confidence level of 0.8 is not representative because of a small sample size in this bin.
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.3459 0.5858
Chain-of-Thought 0.3065 0.6379
Surrogate Token 0.3686 0.5923
Sequence Probability 0.3156 0.7804
Self-Consistency 0.1134 0.8213
Trained Probe 0.0189 0.8173
Rewarding Doubt (ours) 0.0226 0.8592
Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting.
(a) Base model
(b) Fine-tuned by Rewarding Doubt
Figure 4: Calibration curves of the zero-shot base model (Verbalize) and the model fine-tuned by Rewarding Doubt.
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.5319 0.6047
Sequence probability 0.3230 0.5942
Trained probe 0.2119 0.6481
Rewarding doubt (ours) 0.1061 0.7268
Table 2: Comparison of methods on the QAMPARI dataset in the Multiple-Answer setting.
Rewarding Doubt breaks the pattern of overconfidence
Before fine-tuning the base model, we can observe in Figure 5(a) that the model is very overconfident and almost always reports a confidence of eight or higher. This pattern is also observed by Xiong et al. (2024), who hypothesize that this overconfidence comes from the supervised pre-training, where the training data usually expresses high confidence, leading the model to mimic these surface-level patterns instead of developing a real sense of confidence. After fine-tuning the model with our method, as shown in Figure 5(b), it expresses confidence scores distributed over a much wider range, including lower confidence values. This shift indicates that the model is able to overcome the learned pattern of overconfidence and develops an improved sense of confidence.
Rewarding Doubt shows significant generalization capabilities between datasets
To assess the generalization capabilities of our method, we evaluate the model trained on TriviaQA in an out-of-domain setting, using the CommonsenseQA and MedQA datasets. The results are presented in Table 3 and Table 4, respectively. Our approach is significantly better able to generalize than the Trained Probe method, which performs poorly on both datasets and completely fails to generalize to CommonsenseQA. Additionally, our fine-tuned model is significantly better calibrated than the base model on MedQA. However, on CommonsenseQA, the ECE of our approach is slightly worse than that of the base model. This reveals a flaw in the ECE metric. Notably, the ECE does not account for the range of confidence values. For instance, if a model were to consistently predict a confidence of 6 for a dataset with an overall accuracy of 60%, the ECE would be optimal. However, such confidence values would lack practical utility, as they would not enable meaningful discrimination between individual cases. Therefore, a low ECE should not be the sole criterion for evaluating confidence estimation. A model’s ability to discriminate between correct and incorrect answers, as measured by the AUROC score, should also be considered. Consequently, the improvement in the AUROC score compared to the zero-shot verbalization, supports the increased utility of the confidence estimates produced by the fine-tuned model. Overall, these generalization experiments demonstrate that the uncertainty indicators the model learns are not entirely dataset-specific; rather, our method enables the model to develop a more generalized understanding of confidence.
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.2820 0.5425
Trained Probe 0.4819 0.5374
Rewarding doubt (ours) 0.2930 0.6385
Table 3: Comparison of generalization results on the CommonsenseQA dataset trained on the TriviaQA dataset.
Method ECE ($\downarrow$) AUROC ($\uparrow$)
Verbalize 0.4480 0.5075
Trained Probe 0.2099 0.5513
Rewarding doubt (ours) 0.1145 0.6649
Table 4: Comparison of generalization results on the MedQA dataset trained on the TriviaQA dataset.
Rewarding Doubt can generalize between settings
Beyond generalization across datasets, we also assess whether our approach transfers between different experimental settings. When evaluating the model trained in the Single-Answer setting directly on a Multiple-Answer task, it has a considerably better calibration compared to the base model, as seen in Table 5. While it does not achieve the same level of calibration as the model trained directly in the Multiple-Fact setting, this suggests potential benefits in future applications, particulary for improving the calibration of facts in written texts. Training on single facts could already improve the calibration in open-text applications, making it a promising approach for scenarios where structured multi-fact training is impractical.
Training Evaluation ECE ($\downarrow$) AUROC ($\uparrow$)
Base model Single fact 0.5875 0.5787
Single fact Single fact 0.1536 0.7240
Base model Multi fact 0.5319 0.6047
Single fact Multi fact 0.1777 0.6617
Multi fact Multi fact 0.1061 0.7268
Table 5: Comparison of the base and fine-tuned model on the Qampari dataset in different settings.
Rewarding Doubt is more resource efficient
An advantage of our approach, not visible in the metrics, is the inference speed. While it initially takes more time for training, inference is highly efficient, as only a small constant number of tokens has to be generated to estimate confidence. In constrast, methods like Chain-of-Thought and Self-Consistency need to generate significantly more tokens for their reasoning process, increasing the computational costs and inference time by several orders of magnitude. Our approach introduces no additional overhead, does not depend on external resources, and directly provides confidence information through verbalization. Thus, the time required to express confidence with our method is significantly less than that of previous approaches.
(a) Base model
(b) Fine-tuned by Rewarding Doubt
Figure 5: Histograms of predicted confidences of the zero-shot base model (Verbalize) and the model fine-tuned on the TriviaQA dataset.
Rewarding Doubt does not effect answer correctness
Training confidence calibration with our method only targets the uncertainty estimation abilities and does not effect the performance on the task as seen by the accuracy in Table 6. Additionally, our initial hypothesis that the model could learn to generate incorrect answers and label them accordingly with a low confidence score, does not appear to be a problem. As shown in Table 6, there is no significant difference in accuracy, calibration, or AUROC scores with and without an additional reward for correct answers. Given that our method rewards the model on its expressed confidence and not the answer, it is unlikely that the model deliberately generates incorrect responses. Nevertheless, as the inclusion of this additional reward does not negatively impact performance, it can be retained to ensure robust confidence estimation without unintended side effects across different settings.
Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
Base model 0.3459 0.5858 0.6310
Bonus 0.0226 0.8592 0.6309
No bonus 0.0219 0.8562 0.6278
Table 6: Comparison of effect of Rewarding Doubt on accuracy on the TriviaQA dataset.
Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, generalize across tasks, and maintain computational efficiency, making it an effective approach for accurate confidence calibration in language models.
Limitations
During our experiments, we observed that the policy occasionally converges to always predicting a fixed confidence value, likely caused by training instabilities. While this did not occur consistently, it suggests a potential area for further refinement, such as adjusting the reward structure or introducing additional constraints to encourage a more stable learning of a diverse confidence distribution. Additionally, our approach is currently limited to settings, where a binary correctness measure can be established. In the future this can be extended to also include continuous correctness measures such as usually applied for free text generation.
6 Conclusion
In this work, we propose Rewarding Doubt, a novel approach that enables LLMs to express confidence in their answers more accurately using natural language. We leverage reinforcement learning with a reward function that incentivizes diverse and well-calibrated confidence expressions. Fine-tuning with our method significantly improves the model’s ability to convey calibrated confidence, effectively reducing the overconfidence commonly observed in LLMs. This not only enhances the trustworthiness in AI-generated responses but also lays the groundwork for more reliable human-AI collaboration, where models can transparently communicate uncertainty, an essential step toward safer and more accountable AI systems.
Acknowledgements
The authors gratefully acknowledge the financial support by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project ThoraXAI (DIK-2302-0002), and the German Research Foundation (DFG, grant 469106425 - NA 620/51-1).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Prompts
For all the question-answering settings, the model is directly prompted to answer a question without a preceding example or context. For our method the model was prompted to answer the question and additionally provide a verbalized confidence. For the other baselines that do not need a verbalized confidence but infer it indirectly, the model is prompted to only give the correct answer. The specifics for multiple-choice are slightly changed but hold mostly the same meaning. The exact prompts for each method can be seen in Table 7 for open questions and Table 8 for multiple-choice questions. The prompts for each Multi-Answer method can be seen in Table 9. We decided not to give the model a role like "expert" in the system prompt but keep it neutral, as we observed that the role we give the model affects the verbalized confidence.
Method Prompt
Verbalize/ Rewarding Doubt You will get questions. Answer with the correct answer. Additionally provide a confidence between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, of how sure you are the answer is correct. A value close to 0 means you think there is a high probability that the answer is wrong. The closer the value is to 10, the higher you think is the probability that the answer is correct. The output should have the format ’Answer: <answer>, Confidence: <confidence>’ and nothing else.
CoT You will get questions. Read the question, analyze step by step, provide your reasoning steps, answer and your confidence in the response. The confidence should be between 0, 1, 2, 3, 4, 5. A value close to 0 means you think there is a high probability that you could be wrong. The closer the value is to 5, the lower you think is the chance that you could be wrong. The output should have the format ’Thoughts: <reasoning steps>, Answer: <just the answer and nothing else>, Confidence: <confidence>’.
Sequence Probability/ Trained Probe You will get questions. Answer with the correct answer only and nothing else.
[…]
Self-Consistency You will get questions. Read the question, analyze step by step, provide your reasoning steps and answer. The output should have the format ’Thoughts: <reasoning steps>, Answer: <just the answer and nothing else>’
Table 7: The prompts for each method for the open short form questions.
Method Prompt
Verbalize/ Rewarding Doubt You will get test questions with possible options. Answer with the correct option. Additionally provide a confidence between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, of how sure you are the answer is correct. A value close to 0 means you think there is a high probability that the answer is wrong. The closer the value is to 10, the higher you think is the probability that the answer is correct. The output should have the format ’Answer: <answer_index>, Confidence: <confidence>’ and nothing else.
CoT You will get test questions with possible options. Read the question, analyze step by step, provide your reasoningsteps, answer and your confidence in the response. The confidence should be between 0, 1, 2, 3, 4, 5. A value close to 0 means you think there is a high probability that you could be wrong. The closer the value is to 5, the lower you think is the chance that you could be wrong. The output should have the format ’Thoughts: <reasoning steps>, Answer: <answer_index>, Confidence: <confidence>’ and nothing else.
Sequence Probability/ Trained Probe You will get test questions with possible options. Answer with the correct option index only and nothing else.
[…]
Self-Consistency You will get test questions with possible options. Read the question, analyze step by step, provide your reasoningsteps and the correct option index. The output should have the format ’Thoughts: <reasoning steps>, Answer: <answer_index>’ and nothing else.
Table 8: The prompts for each method for multiple choice questions.
Method Prompt
[…]
Table 9: The prompts for each method for multiple fact questions.

### v3

@@
-Rewarding Doubt: A Reinforcement Learning Approach to Confidence Calibration of Large Language Models
+Rewarding Doubt: A Reinforcement Learning Approach to Calibrated Confidence Expression of Large Language Models
 Abstract
 A safe and trustworthy use of Large Language Models (LLMs) requires an accurate expression of confidence in their answers.
-We introduce a novel Reinforcement Learning (RL) approach for LLM calibration that fine-tunes LLMs to elicit calibrated confidence estimations in their answers to factual questions.
-We model the problem as a betting game where the model predicts a confidence score together with every answer, and design a reward function that penalizes both over and under-confidence.
-We prove that under our reward design an optimal policy would result in a perfectly calibrated confidence estimation.
-Our experiments demonstrate significantly improved confidence calibration and generalization to new tasks without re-training, indicating that our approach teaches a general confidence awareness.
-This approach enables the training of inherently calibrated LLMs.
-Machine Learning, ICML
+We propose a novel Reinforcement Learning approach that allows to directly fine-tune LLMs to express calibrated confidence estimates alongside their answers to factual questions.
+Our method optimizes a reward based on the logarithmic scoring rule, explicitly penalizing both over- and under-confidence.
+This encourages the model to align its confidence estimates with the actual predictive accuracy.
+The optimal policy under our reward design would result in perfectly calibrated confidence expressions.
+Unlike prior approaches that decouple confidence estimation from response generation, our method integrates confidence calibration seamlessly into the generative process of the LLM.
+Empirically, we demonstrate that models trained with our approach exhibit substantially improved calibration and generalize to unseen tasks without further fine-tuning, suggesting the emergence of general confidence awareness.
+We provide our training and evaluation code in the supplementary and will make it publicly available upon acceptance.
 1 Introduction
-In human intelligence and inter-human interaction, the ability to understand our own uncertainty and communicate our doubts to others is fundamental for effective decision-making, collaboration, and learning (Cosmides & Tooby, 1996; Xiong et al., 2024).
+In human intelligence and inter-human interaction, the ability to understand our own uncertainty and communicate our doubts to others is fundamental for effective decision-making, collaboration, and learning [7, 37].
 Similarly, for Large Language Models (LLMs) to be safely used in real-world applications, especially when humans and AI systems work together, they must not only generate accurate information but also communicate their confidence in that information.
-While LLMs have demonstrated impressive capabilities in natural language understanding, question answering and text summarization (Touvron et al., 2023; Chiang et al., 2023; Achiam et al., 2023), LLMs still face significant limitations, such as their tendency to generate inaccurate information, often referred to as hallucinations (Hadi et al., 2023).
+While LLMs have demonstrated impressive capabilities in natural language understanding, question answering and text summarization [32, 6, 1], LLMs still face significant limitations, such as their tendency to generate inaccurate information, often referred to as hallucinations [11].
 This raises concerns about their reliability, particularly in real-world applications where trustworthiness is essential.
-Especially in high-stakes environments such as medical diagnosis, where LLMs are starting to become support tools for professionals (Moor et al., 2023; Pellegrini et al., 2023; Tu et al., 2024; Bani-Harouni et al., 2024), overconfident predictions including factual errors or hallucinations could have serious consequences for patient health.
-Also, in customer service or legal consultation (Shi et al., 2024; Sun et al., 2024), LLMs need to express uncertainty and defer complex queries to human representatives when unsure to avoid misinformed decisions.
-Reliable confidence estimation would enable these systems to flag uncertain outputs for human review, ensuring that crucial decisions are not made based on uncertain LLM outputs.
+Especially in high-stakes environments such as medical diagnosis, where LLMs are starting to become support tools for professionals [23, 25, 33, 4], overconfident predictions including factual errors or hallucinations could have serious consequences for patient health.
+Also, in customer service or legal consultation [26, 28], LLMs need to express uncertainty and defer complex queries to human representatives when unsure to avoid misinformed decisions.
+Reliable confidence estimation and expression would enable these systems to flag uncertain outputs for human review, ensuring that crucial decisions are not made based on uncertain LLM outputs.
 To allow risk estimation while using LLM-generated output, model confidence should be calibrated, meaning that the expressed numerical confidence should be equal to the probability of the model’s answer being correct.
@@
 Our reward function is designed so that the best reward is given when the confidence and the epistemic probability of being correct are the same, thus incentivizing the expression of calibrated confidences.
-Many previous methods for confidence estimation lack in calibration performance as they do not train the model and instead infer the confidence from the internal state in a zero-shot setup (Huang et al., 2023b; Kuhn et al., 2023; Duan et al., 2024).
+Many previous methods for confidence estimation lack in calibration performance as they do not train the model and instead infer the confidence from the internal state in a zero-shot setup [14, 19, 8].
 Additionally, this does not give models an inherent awareness of confidence.
-Other trained methods in this area generally do not explicitly optimize for calibration but instead for a binary correctness target (Azaria & Mitchell, 2023; Kapoor et al., 2024).
-While this ensures strong performance in terms of discrimination, it overlooks the practical importance of fine-grained and well-calibrated confidence scores.
-Without proper calibration, users cannot reliably interpret confidence values, making it difficult to determine when to trust a model’s predictions.
-This is particularly problematic in real-world applications where decision thresholds vary based on context.
-For example, a high-stakes clinical decision requires high confidence, while a low-stakes trivia question might tolerate lower confidence.
-By contrast, well-calibrated models enable more flexible and trustworthy human interaction, allowing users to dynamically adjust their confidence requirements based on the stakes of a given decision.
+Other trained methods in this area decouple the uncertainty estimation from the text generation process [3, 18].
+This approach optimizes for calibrated confidence estimation but does not enable the uncertainty-awareness and expression in the model itself.
 Targeting these limitations, we propose a novel reinforcement-learning (RL) approach for teaching LLMs to express their calibrated confidence, encouraging a granular, accurate estimation of the confidence level in the training objective.
 For this, we model confidence estimation as a betting game: a high-confidence answer would warrant a larger bet, reflecting a strong belief in its correctness, while a lower confidence score would suggest caution.
-Our reward function is designed to reflect the risk-reward tradeoff in betting, as illustrated in Figure 1.
-It increases the reward when a correct answer is given with high confidence, simulating the higher stakes of big bets.
+Central to our method is a reward function based on the logarithmic scoring rule, a strictly proper scoring rule.
+We are the first to optimize this function through reinforcement-learning-based policy optimization, leveraging its calibration properties for directly and seamlessly training confidence calibration in LLM generations.
+This reward function captures the fundamental risk-reward tradeoff in probabilistic decision-making, as illustrated in Figure 1.
+It increases the reward when a correct answer is given with high confidence, simulating the higher potential return of big bets.
 Conversely, it penalizes incorrect answers more when they are made with high confidence, discouraging overconfidence.
 This ensures that both uncertainty and confidence are appropriately factored into the reward system.
-With our reward design, the model learns to align its predicted confidence with the accuracy of its output, encouraging granular and calibrated confidence scoring, as a calibrated confidence estimation will provably result in the highest reward during training.
+As a proper scoring rule, optimizing the reward function trains the model to align its predicted confidence with the accuracy of its output, encouraging granular and calibrated confidence scoring.
+A calibrated confidence estimation will provably result in the highest reward during training.
 This not only improves the trustworthiness of LLMs in collaborative human-AI scenarios but also helps users better assess when AI tools should be trusted, double-checked, or deferred to human expertise.
-Figure 2: Overview of our reinforcement learning framework: The LLM is prompted to answer a question and provide the confidence in this answer.
-The answer is checked for correctness by a judge function and the reward is calculated based on the correctness and the confidence.
-Correct answers with high confidences are rewarded highly, but also penalized heavily when incorrect.
 2 Related Works
-Confidence estimation and calibration have a long history in machine learning and natural language processing (Wang, 2024).
+Confidence estimation and calibration have a long history in machine learning and natural language processing [34].
 With the rise of LLMs, a lot of attention has moved to extracting and improving confidence estimation from these new models.
-There is a wide range of methods that can roughly be divided into black-box and white-box methods (Geng et al., 2024).
+There is a wide range of methods that can roughly be divided into black-box and white-box methods [9].
 Black-box methods require only the model generations for uncertainty estimation and do not depend on any knowledge of the model’s internal state.
@@
 Xiong et al.
-(2024) showed that prompting techniques like chain-of-thought (CoT) (Wei et al., 2022) can, to a limited extend, help the model in the calibration of its confidence expression.
-Based on the assumption that uncertainty in its generation would prime the model to hallucinate information, consistency-based methods (Manakul et al., 2023; Wang et al., 2022) measure the similarity between multiple generations and show that high variance in the generation correlates with high uncertainty.
+[37] showed that prompting techniques like chain-of-thought (CoT) [36] can, to a limited extend, help the model in the calibration of its confidence expression.
+Based on the assumption that uncertainty in its generation would prime the model to hallucinate information, consistency-based methods [21, 35] measure the similarity between multiple generations and show that high variance in the generation correlates with high uncertainty.
 Wang et al.
-(2022) propose using CoT to produce various reasoning pathways towards an answer.
+[35] propose using CoT to produce various reasoning pathways towards an answer.
 By marginalizing out the reasoning pathways, they provide the most consistent answer as the one most probably correct and offer the consistency measure as an estimation of the model’s confidence.
@@
 Huang et al.
-(2023a) calculate the mean probability of the generated tokens with the intuition that a model that generates an answer with high probability tokens also has a high confidence in this answer.
+[14] calculate the mean probability of the generated tokens with the intuition that a model that generates an answer with high probability tokens also has a high confidence in this answer.
 Kuhn et al.
-(2023) and (Duan et al., 2024) expand on this idea by looking at the entropy of the generated tokens while also taking the semantic relevance into account.
+[19] and [8] expand on this idea by looking at the entropy of the generated tokens while also taking the semantic relevance into account.
 Kadavath et al.
-(2022) showed that LLMs can self-evaluate the correctness of their own generation.
+[17] showed that LLMs can self-evaluate the correctness of their own generation.
 They prompt the model to provide an answer followed by a judgement whether its own answer is "true" or "false".
@@
 Xiong et al.
-(2024) propose a method that trains a linear probe for assessing the model’s confidence based on the activations of the LLM’s hidden state.
+[37] propose a method that trains a linear probe for assessing the model’s confidence based on the activations of the LLM’s hidden state.
 They evaluate the application of the probe to various layers and train it with correctness labels of the task.
@@
 Building on this approach, Mielke et al.
-(2022) give an LLM access to the confidence extracted by the external probe and fine-tune it to express this confidence.
+[22] give an LLM access to the confidence extracted by the external probe and fine-tune it to express this confidence.
 However, the calibration performance is upper-bounded by the ability of the trained probe to assess the model’s confidence accurately.
-Fine-tuning LLMs using Reinforcement Learning from human feedback (RLHF) has received increasing attention for aligning the models with human preferences, avoiding toxicity or the generation of harmful content (Ouyang et al., 2022; Lu et al., 2022).
-Outside of RLHF, RL has been used to fine-tune LLMs for agentic interaction in textual environments (Zhou et al., 2023; Carta et al., 2023).
+Stengel-Eskin et al.
+[27] propose a DPO-based method, LACIE, simulating an interaction between a generation and a listener model, where the speaker model generates answers, and the listener model evaluates them based on perceived confidence.
+By training with Direct Preference Optimization (DPO) on data from TriviaQA, LACIE rewards accurate confidence expression, including correct answers accepted by the listener model and incorrect, rejected ones, while penalizing misleading confidences.
+The confidence expression is calibrated to how the listener model interprets confidence cues, depending on its ability to judge confidence and aligning perceived confidence rather than a fact-based numerical confidence score.
+However, none of these methods provide theoretical guarantees or provable optimality regarding the trained calibration of the expressed confidence.
+Simultaneously, fine-tuning LLMs using Reinforcement Learning from human feedback (RLHF) has received increasing attention for aligning the models with human preferences, avoiding toxicity or the generation of harmful content [24, 20].
+Outside of RLHF, RL has been used to fine-tune LLMs for agentic interaction in textual environments [40, 5].
 Confidence estimation of LLMs using RL is a novel research direction that, to the best of our knowledge, has only been approached by Tao et al.
-(2024) and Xu et al.
-(2024).
+[30] and Xu et al.
+[38].
 Tao et al.
-(2024) build on the Reinforcement Learning with Human Feedback (RLHF) paradigm and design a reward function that aligns the verbalized confidence of the model with the preference rating predicted by the RLHF reward model.
+[30] build on the Reinforcement Learning with Human Feedback (RLHF) paradigm and design a reward function that aligns the verbalized confidence of the model with the preference rating predicted by the RLHF reward model.
 However, this approach can not evaluate model confidence in factual statements and depends on human-annotated preference data.
 Xu et al.
-(2024) introduce a novel Reinforcement Learning from Knowledge Feedback (RLKF) framework, which aims to teach a model to refuse to answer questions outside their knowledge scope, reducing hallucinations.
-However, this does not quantify confidence when giving an answer.
+[38] introduce a novel Reinforcement Learning from Knowledge Feedback (RLKF) framework, which aims to teach a model to refuse to answer questions outside their knowledge scope, reducing hallucinations, but does not quantify confidence when giving an answer.
+In contrast, our method directly optimizes for factual calibration using a theoretically grounded, proper scoring rule as the reward signal, enabling the model to develop intrinsic uncertainty awareness without requiring external preference models, knowledge supervision, or post-hoc calibration techniques, while at the same time seamlessly integrating calibrated confidence expression into the LLMs response generation.
 3 Rewarding Doubt
@@
 Based on the correctness of the answer, and the expressed confidence, we reward the model, incentivizing it to express a calibrated confidence.
-Formally, let the model be provided with a textual question or request $q$, prompting an answer-confidence pair $(a,\hat{p})$ as response, where $a$ is a textual answer with binary correctness value, and $0\leq\hat{p}\leq 1$ is a numerical confidence score representing the subjective probability the model assigns to answer $a$ being correct.
+Figure 2: Overview of our reinforcement learning framework: The LLM is prompted to answer a question and provide the confidence in this answer.
+The answer is checked for correctness by a judge function and the reward is calculated based on the correctness and the confidence.
+Correct answers with high confidences are rewarded highly, but also penalized heavily when incorrect.
+Formally, let the model be provided with a textual question or request $q$, resulting in an answer-confidence pair $(a,\hat{p})$ as response, where $a$ is a textual answer with binary correctness value, and $0\leq\hat{p}\leq 1$ is a numerical confidence score representing the subjective probability the model assigns to answer $a$ being correct.
 We train this subjective probability assessment to align with the true epistemic probability $p^{*}$, which represents the actual likelihood of correctness given the model’s internal knowledge state.
@@
 We design our reward as a logarithmic scoring function, fulfilling the described optimality conditions:
+$R(a,\hat{p},j)=\begin{cases}log(\max(\hat{p},\epsilon)),&\text{if $j(a)$ = 1 (correct)}\\ log(\min(1-\hat{p},1-\epsilon)),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
+where $\epsilon>0$ is a small positive constant introduced for numerical stability to avoid evaluating the logarithm at zero.
+The normalized reward for correct and incorrect answers for each confidence is visualized in Figure 3.
+Proposition 1 (Optimality implies Calibration).
+The expected reward $\mathbb{E}[R(a,\hat{p},j)]$ is maximized for each sample when $\hat{p}=p^{*}$ and the optimal policy under the reward design is thus perfectly calibrated.
+The proof of Proposition 1 is analogous to the proof that the logarithmic scoring rule is a proper scoring rule.
+We provide it in full in Appendix B and discuss the influence of the clipping on the optimality of the reward function.
+Figure 3: The rewards for each confidence value for correct and incorrect answers.
+The closer the confidence is to ten or zero, respectively, the higher is the reward.
+At the same time, the possible punishment increases to a greater extent.
+The model has to learn when the trade-off between those two possibilities is worthwhile.
+4 Experimental Setup
+We evaluate our method in both Single-Answer and Multiple-Answer settings.
+We prompt the model to provide a confidence for each answer as an integer between 0 and 10, which we normalize for the reward calculation.
+A confidence of zero is defined as the model being certain that the answer is incorrect, while ten is defined as the model being certain the answer is correct.
+We normalize the reward function to the range of $[-1,1]$.
+In the Single-Answer setting we train the model on the TriviaQA dataset [16], which contains question-answer-evidence triplets, from which we only use the questions and answers.
+For generalization experiments, we evaluate our method on CommonsenseQA [29] and MedQA [15], which are multiple-choice question datasets in the commonsense and medical domain, respectively.
+For the Multiple-Answer setting, we train on the QAMPARI dataset [2], which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
+In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought [37], Top-K [31], Surrogate Token [17], Sequence Probability [14] and Self-Consistency [35] as zero-shot methods, LACIE [27], which uses DPO for optimizing confidence expression and Trained Probe [3], which employs supervised training of an external probe for estimation model.
+We also compare to the non-finetuned base model in a zero-shot manner, using the same prompt as our Rewarding Doubt method and refer to this setup as Verbalize.
+In the Multiple-Answer setting we compare to Trained Probe and Sequence probability, as those methods are the best performing zero-shot and trained baselines in the Single-Answer setting.
+LACIE does not report results for this dataset, thus we can only compare on TriviaQA.
+We report our results using the Expected Calibration Error (ECE) and the Area Under the Receiver Operating Characteristic Curve (AUROC) metric.
+The ECE measures the difference between a model’s predicted confidence and its actual accuracy, assessing how well confidences align with real outcomes.
+AUROC assesses a model’s ability to distinguish between classes, with a score of 0.5 indicating random guessing and 1.0 indicating perfect classification.
+Additionally, we visualize the calibration with calibration curves, where a well-calibrated model lies close to the 45° line and large deviations show a high miscalibration.
+Response Generation
+To calibrate and reward the model only on the confidences and not the answers we separate generation in two steps during training: Answer and confidence generation.
+Answers are generated first and afterwards treated as fixed inputs alongside the question, while the confidence is generated in a separate generation step and considered as sole target for optimization.
+Like this, we ensure that answer generation is disentangled from the optimization process, ensuring the answer correctness is not affected by our confidence calibration training.
+Correctness Assessment
+For the multiple-choice datasets MedQA and CommonsenseQA, we evaluate correctness using the exact string matching between the model’s response and the ground truth answer.
+For the TriviaQA and QAMPARI datasets, we use the F1 score of word overlap to measure the similarity between the model’s response and the ground truth candidates.
+The F1 score is calculated for each candidate and the maximum score is considered the final score.
+We consider an answer as correct if its score exceeds a threshold of 0.5.
+5 Implementation Details
+Training
+As base model for our experiments we use Meta-Llama-3-8B-Instruct [10].
+We employ the 4-bit quantized performance-optimized model version by Unsloth AI [12].
+Additionally, we apply LoRA fine-tuning [13], reducing the amount of trainable parameters from 8,030,265,345 to 41,947,137.
+For the Single-Answer setting we train the model for two epochs with a learning rate of 1e-5.
+For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 24,000 steps with a batchsize of eight and a learning rate of 1e-5 with a reward scale of five.
+All models are trained on one Nvidia A40 with each training run taking seven days.
+On average the model generated approximately 3.4 answers per fact.
+If the model fails to generate an answer in the specified format, it is penalized with an out-of-format reward of -3.
+Baselines
+For the Sequence Probability, we compute the average probability for each token in the response.
+In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric [39].
+For the trained probe [3], the original study introduced a custom dataset comprising short statements classified as either true or false.
+The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements.
+To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers.
+The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework.
+For the architecture of the MLP, we employed the same design as Azaria and Mitchell [3] and train it for four epochs with a learning rate of 1e-4 until convergence.
+The exact prompts used for each baseline are provided in Appendix A.
+6 Results and Discussion
+Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting. *Results are from the original paper [27].
+Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
+Verbalize 0.3459 0.5858 0.6310
+Chain-of-Thought 0.3065 0.6379 0.6273
+Top-K 0.1611 0.6673 0.6023
+Surrogate Token 0.3686 0.5923 0.5933
+Sequence Probability 0.3156 0.7804 0.5955
+Self-Consistency 0.1134 0.8213 0.6224
+LACIE* 0.1200 0.7200 n/a
+Trained Probe 0.0189 0.8173 0.5925
+Rewarding Doubt (ours) 0.0226 0.8592 0.6309
+This section presents and discusses the key findings of our experiments for both Single and Multiple-Answer tasks and the generalization to out-of-domain datasets.
+Calibration Performance
+To assess how well our approach improves calibration, we compare it against the zero-shot LLM baseline (Verbalize) and several established methods in both Single-Answer and Multiple-Answer question-answering tasks.
+Results for the Single-Answer setting on TriviaQA are presented in Table 1, and those for the Multiple-Answer setting on QAMPARI appear in Table 2.
+Across both tasks, Rewarding Doubt substantially improves the model’s confidence calibration over zero-shot verbalization.
+In the Single-Answer setting on TriviaQA, Rewarding Doubt achieves an ECE of 0.0226 and an AUROC of 0.8592, clearly outperforming all zero-shot baselines as well as LACIE, which is based on DPO-based optimization.
+The second fine-tuned method, Trained Probe, which relies on supervised fine-tuning, reports a slightly lower ECE (0.0189), both methods achieve near-perfect results.
+Further the AUROC of Rewarding Doubt is notably higher, suggesting that although both methods offer strong calibration, Rewarding Doubt better discriminates between correct and incorrect answers.
+In the Multiple-Answer setting on QAMPARI, Rewarding Doubt also outperforms baselines, achieving an ECE of 0.0816 and an AUROC of 0.6947.
+In comparison, Verbalize, Sequence Probability, and Trained Probe perform notably worse.
+Our findings support the claim by Azaria and Mitchell [3] that a model’s internal state encodes information about the truthfulness of statements, which can serve as an indicator of uncertainty.
+However, without fine-tuning, the model struggles to utilize this internal information effectively.
+Our approach enables the model to make use of this correlation and translate it into an accurate expression of the probability that a given answer is correct.
+The calibration curves in Figure 4 further illustrate these improvements.
+For both TriviaQA and QAMPARI, the fine-tuned model’s confidence much more closely aligns with the ideal 45° line than the zero-shot Verbalize baseline.
+Additionally, we observe a shift in the confidence distribution after fine-tuning.
+As shown in Figure 5, in a zero-shot setting the LLM (Verbalize) predominantly assigns high confidence scores (8 or above), reflecting overconfidence, a pattern also noted by Xiong et al.
+[37], who attribute it to supervised pretraining that favors confident expressions.
+After fine-tuning with Rewarding Doubt, the model’s confidence scores (shown in Figure 5(b)) span a wider range, including lower values, indicating a more nuanced expression of uncertainty.
+This shift suggests that fine-tuning mitigates overconfidence and better aligns the model’s confidence with its actual performance.
+(a) Base model
+(b) Fine-tuned by Rewarding Doubt
+Figure 4: Calibration curves of the zero-shot base model (Verbalize) and the model fine-tuned by Rewarding Doubt.
+Table 2: Comparison of methods on the QAMPARI dataset in the Multiple-Answer setting.
+Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
+Verbalize 0.5319 0.6047 0.2549
+Sequence probability 0.5324 0.5942 0.1927
+Trained probe 0.1117 0.6481 0.2233
+Rewarding doubt (ours) 0.0816 0.6947 0.2480
+(a) Base model
+(b) Fine-tuned by Rewarding Doubt
+Figure 5: Histograms of predicted confidences of the zero-shot base model (Verbalize) and the model fine-tuned on the TriviaQA dataset.
+Table 3: Comparison of generalization results on the CommonsenseQA dataset trained on the TriviaQA dataset.
+Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
+Verbalize 0.2820 0.5425 0.6860
+Trained Probe 0.4819 0.5374 0.7108
+Rewarding doubt (ours) 0.2930 0.6385 0.7163
+Stability of Answer Correctness
+Training confidence calibration with our method only targets the uncertainty estimation abilities and does not aim to alter the responses of the model.
+This is achieved by only rewarding the model on its expressed confidence, while the answer is generated beforehand independently from the model update step.
+Our results show a stable accuracy for all experiments without notable differences in accuracy between the base model (Verbalize) and the model adapted with Rewarding Doubt, showing that confidence calibration training with Rewarding Doubt does not affect task performance.
+Generalization Capabilities
+To assess the generalization abilities of Rewarding Doubt, we evaluated the model trained on TriviaQA in out-of-domain settings using the CommonsenseQA [29] and MedQA [15] datasets.
+Results are shown in Table 3 and Table 4, respectively.
+On MedQA, Rewarding Doubt significantly outperforms Verbalize in both metrics, while on CommonsenseQA, it achieves a comparable ECE, however paired with a much higher AUROC.
+This discrepancy highlights a limitation of relying solely on ECE for evaluating calibration.
+ECE does not reflect how well a model discriminates between correct and incorrect predictions across different confidence levels.
+A model consistently assigning moderate confidence values could appear well-calibrated under ECE, yet fail to offer meaningful distinctions between uncertain and certain cases.
+The AUROC, by contrast, directly measures this discriminative ability.
+Thus, the substantial improvements in AUROC underscore that Rewarding Doubt produces more useful and actionable confidence estimates.
+Compared to the Trained Probe baseline, Rewarding Doubt consistently outperforms with significantly improved metrics, showing a stronger ability to generalize to new datasets.
+Table 4: Comparison of generalization results on the MedQA dataset trained on the TriviaQA dataset.
+Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
+Verbalize 0.4480 0.5075 0.5066
+Trained Probe 0.2099 0.5513 0.5051
+Rewarding doubt (ours) 0.1145 0.6649 0.5161
+We also explore generalization across experimental settings by applying a model trained in a Single-Answer setting to a Multiple-Answer task.
+Although it does not reach the performance of a model trained specifically for that task, it still outperforms baseline methods considerably, demonstrating transferability of the learning confidence estimation patterns.
+This suggests promising applications for improving confidence estimation in more complex or less structured scenarios, such as fact verification and calibration in free-text generation, even when specialized training data is unavailable.
+Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, while generalizing across tasks, and maintaining stable task performance, making it an effective approach for accurate confidence calibration and expression in language models.
+Beyond improvements in calibration quality, our method also offers practical advantages.
+While fine-tuning requires an initial training investment, inference remains highly efficient, as only a small, constant number of tokens need to be generated to express confidence.
+In contrast, zero-shot methods like Chain-of-Thought and Self-Consistency have substantial computational overhead during inference by requiring lengthy reasoning chains or multiple generations.
+Rewarding Doubt introduces no such overhead, does not rely on an additional model, and directly provides actionable confidence estimates through simple verbalization directly by the LLM, making it highly suitable for real-world deployment.
+Table 5: Comparison of the base and fine-tuned model on the Qampari dataset in different settings.
+Training Evaluation ECE ($\downarrow$) AUROC ($\uparrow$)
+Base model Single fact 0.5875 0.5787
+Single fact Single fact 0.1536 0.7240
+Base model Multi fact 0.5319 0.6047
+Single fact Multi fact 0.1777 0.6617
+Multi fact Multi fact 0.1061 0.7268
+Limitations
+During our experiments, we observed that the policy occasionally converges to always predicting a fixed confidence value, likely caused by training instabilities.
+While this did not occur consistently, it suggests a potential area for further refinement, such as adjusting the reward structure or introducing additional constraints to encourage a more stable learning of a diverse confidence distribution.
+Additionally, our approach is currently limited to settings, where a binary correctness measure can be established.
+In the future this can be extended to also include continuous correctness measures such as usually applied for free text generation.
+7 Conclusion
+In this work, we propose Rewarding Doubt, a novel approach that enables LLMs to express confidence in their answers more accurately using natural language.
+We leverage reinforcement learning with a reward function based on the logarithmic scoring rule that incentivizes well-calibrated confidence expressions.
+Fine-tuning with our method significantly improves the model’s ability to estimate a calibrated confidence, effectively reducing the overconfidence patterns commonly observed in LLMs.
+This not only enhances the trustworthiness in AI-generated responses but also lays the groundwork for more reliable human-AI collaboration, where models can transparently communicate uncertainty, an essential step toward safer and more accountable AI systems.
+Acknowledgements
+The authors gratefully acknowledge the financial support by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project ThoraXAI (DIK-2302-0002), and the German Research Foundation (DFG, grant 469106425 - NA 620/51-1).
+Appendix A Prompts
+For all the question-answering settings, the model is directly prompted to answer a question without a preceding example or context.
+For our method the model was prompted to answer the question and additionally provide a verbalized confidence.
+For the other baselines that do not need a verbalized confidence but infer it indirectly, the model is prompted to only give the correct answer.
+The specifics for multiple-choice are slightly changed but hold mostly the same meaning.
+The exact prompts for each method can be seen in Table 6 for open questions and Table 7 for multiple-choice questions.
+The prompts for each Multi-Answer method can be seen in Table 8.
+We decided not to give the model a role like "expert" in the system prompt but keep it neutral, as we observed that the role we give the model affects the verbalized confidence.
+Table 6: The prompts for each method for the open short form questions.
+Method Prompt
+Verbalize/ Rewarding Doubt You will get questions.
+Answer with the correct answer.
+Additionally provide a confidence between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, of how sure you are the answer is correct.
+A value close to 0 means you think there is a high probability that the answer is wrong.
+The closer the value is to 10, the higher you think is the probability that the answer is correct.
+The output should have the format ’Answer: <answer>, Confidence: <confidence>’ and nothing else.
+CoT You will get questions.
+Read the question, analyze step by step, provide your reasoning steps, answer and your confidence in the response.
+The confidence should be between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10.
+A value close to 0 means you think there is a high probability that you could be wrong.
+The closer the value is to 5, the lower you think is the chance that you could be wrong.
+The output should have the format ’Thoughts: <reasoning steps>, Answer: <just the answer and nothing else>, Confidence: <confidence>’.
+[…]
+Sequence Probability/ Trained Probe You will get questions. Answer with the correct answer only and nothing else.
+[…]
+Self-Consistency You will get questions.
+Read the question, analyze step by step, provide your reasoning steps and answer.
+The output should have the format ’Thoughts: <reasoning steps>, Answer: <just the answer and nothing else>’
+Table 7: The prompts for each method for multiple choice questions.
+Method Prompt
+Verbalize/ Rewarding Doubt You will get test questions with possible options.
+Answer with the correct option.
+Additionally provide a confidence between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, of how sure you are the answer is correct.
+A value close to 0 means you think there is a high probability that the answer is wrong.
+The closer the value is to 10, the higher you think is the probability that the answer is correct.
+The output should have the format ’Answer: <answer_index>, Confidence: <confidence>’ and nothing else.
+CoT You will get test questions with possible options.
+Read the question, analyze step by step, provide your reasoningsteps, answer and your confidence in the response.
+The confidence should be between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10.
+A value close to 0 means you think there is a high probability that you could be wrong.
+The closer the value is to 5, the lower you think is the chance that you could be wrong.
+The output should have the format ’Thoughts: <reasoning steps>, Answer: <answer_index>, Confidence: <confidence>’ and nothing else.
+Sequence Probability/ Trained Probe You will get test questions with possible options. Answer with the correct option index only and nothing else.
+[…]
+Self-Consistency You will get test questions with possible options.
+Read the question, analyze step by step, provide your reasoningsteps and the correct option index.
+The output should have the format ’Thoughts: <reasoning steps>, Answer: <answer_index>’ and nothing else.
+Table 8: The prompts for each method for multiple fact questions.
+Method Prompt
+[…]
+[…]
+Appendix B Proof
+In the following, we prove Proposition 1 with the reward function
 $R(a,\hat{p},j)=\begin{cases}log(\hat{p}),&\text{if $j(a)$ = 1 (correct)}\\ log(1-\hat{p}),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
-In the following, we prove that $R(a,\hat{p},j)$ is maximized when $\hat{p}=p^{*}$ and thus optimizes for a perfect calibration:
 Proof.
-Let $f(\hat{p})$ be the expected reward for all values of $\hat{p}$ and $p^{*}$:
+The proof is analogous to the proof that the logarithmic scoring function is a proper scoring function.
+Let $f(\hat{p})=\mathbb{E}[R(a,\hat{p},j)]$ be the expected reward for all values of $\hat{p}$ and $p^{*}$:
 $f(\hat{p})=p^{*}\log(\hat{p})\;+\;(1-p^{*})\log\bigl{(}1-\hat{p}\bigr{)}.$
@@
 is strictly negative for $\hat{p}\in(0,1)$. Hence, $f(\hat{p})$ is concave and has its global maximum at $\hat{p}=p^{*}$. ∎
-Since the logarithm of zero is undefined, we set the actual lower limit of the confidence $\hat{p}$ to 0.001 and the upper limit to 0.999.
-The normalized reward for correct and incorrect answers for each confidence is visualized in Figure 3.
-Our reward function rewards incorrect responses with a confidence of zero equally to correct responses with a maximum confidence.
-That means the model could exploit the reward function by always responding with a wrong answer and a confidence of zero.
-To counteract this possible scenario, we add a small constant value to the reward function if the answer is correct.
-This does not change the location of the maximum of the reward function and thus does not alter the optimality but discourages the model from systematically responding incorrectly.
-Figure 3: The rewards for each confidence value for correct and incorrect answers.
-The closer the confidence is to ten or zero, respectively, the higher is the reward.
-At the same time, the possible punishment increases to a greater extent.
-The model has to learn when the trade-off between those two possibilities is worthwhile.
-4 Experimental Setup
-We evaluate our method in both Single-Answer and Multiple-Answer settings.
-We prompt the model to provide a confidence for each answer as an integer between 0 and 10, which we normalize for the reward calculation.
-A confidence of 0 is defined as the model being certain that the answer is incorrect, while 10 is defined as the model being certain the answer is correct.
-We normalize the reward function to the range of $[-1,1]$.
-In the Single-Answer setting we train the model on the TriviaQA dataset (Joshi et al., 2017), which contains question-answer-evidence triplets, from which we only use the questions and answers.
-For generalization experiments, we evaluate our method on CommonsenseQA (Talmor et al., 2019) and MedQA (Jin et al., 2020), which are multiple-choice question datasets in the commonsense and medical domain, respectively.
-For the Multiple-Answer setting, we train on the QAMPARI dataset (Amouyal et al., 2023), which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
-In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought (Xiong et al., 2024), Surrogate Token (Kadavath et al., 2022), Sequence Probability (Huang et al., 2023b), Self-Consistency (Wang et al., 2022) and Trained Probe (Azaria & Mitchell, 2023).
-We also compare to the non-finetuned base model in a zero-shot manner, using the same prompt as our Rewarding Doubt method and refer to this setup as Verbalize.
-In the Multiple-Answer setting we compare only to Trained Probe and Sequence probability, as Chain-of-Tought and Self-Consistency have a high runtime in the Multiple-Answer setting and under-performed in the Single-Answer setting.
-We report our results using the Expected Calibration Error (ECE) and the Area Under the Receiver Operating Characteristic Curve (AUROC) metric.
-The ECE measures the difference between a model’s predicted confidence and its actual accuracy, assessing how well confidences align with real outcomes.
-AUROC assesses a model’s ability to distinguish between classes, with a score of 0.5 indicating random guessing and 1.0 indicating perfect classification.
-Additionally, we visualize the calibration with calibration curves, where a well-calibrated model lies close to the 45° line and large deviations show a high miscalibration.
-4.1 Response Generation
-To calibrate and reward the model only on the confidences and not the answers we separate generation in two steps.
-In the Single-Answer-Setting the model is prompted to generate responses in the format "Answer: <answer>, Confidence: <confidence>".
-In the first generation step an additional stopping criterion is the "Confidence" token, ensuring the model stops before generating the confidence.
-This response is then used as input to generate the confidence.
-For multiple answers, the model is prompted to list all its answers together with a confidence score, where each answer-confidence pair has the same format as in the Single-Answer setting.
-The model first generates all answers and their corresponding confidences.
-Individual answers and confidences are parsed with regular expressions, and new queries to generate a confidence for each answer are created by appending tokens preceding the previous confidence token.
-4.2 Correctness Assessment
-For the multiple-choice datasets MedQA and CommonsenseQA, we evaluate correctness using the exact string matching between the model’s response and the ground truth answer.
-For the TriviaQA and QAMPARI datasets, we use the F1 score of word overlap to measure the similarity between the model’s response and the ground truth candidates.
-The F1 score is calculated for each candidate and the maximum score is considered the final score.
-We consider an answer as correct if its score exceeds a threshold of 0.5.
-4.3 Implementation Details
-4.3.1 Training
-As base model for our experiments we use Meta-Llama-3-8B-Instruct (Grattafiori et al., 2024).
-We employ the 4-bit quantized performance-optimized model version by Unsloth AI (Han et al., 2023).
-Additionally, we apply LoRA fine-tuning (Hu et al., 2022), reducing the amount of trainable parameters from 8,030,265,345 to 41,947,137.
-For the Single-Answer setting we train the model for two epochs with a learning rate of 1e-5.
-For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 32,000 steps with a batchsize of eight and a learning rate of 1e-5 with a reward scale of five.
-On average the model generated approximately 3.4 answers per fact.
-If the model fails to generate an answer in the specified format, it is penalized with an out-of-format reward of -3.
-For a correct answer we additionally reward the model with 0.25.
-4.3.2 Baselines
-For the Sequence Probability, we compute the average probability for each token in the response.
-In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric (Zhang et al., 2019).
-For the trained probe (Azaria & Mitchell, 2023), the original study introduced a custom dataset comprising short statements classified as either true or false.
-The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements.
-To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers.
-The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework.
-For the architecture of the MLP, we employed the same design as Azaria & Mitchell (2023) and train it for four epochs with a learning rate of 1e-4 until convergence.
-The exact prompts used for each baseline are provided Appendix A.
-5 Results and Discussion
-This section presents and discusses the key findings of our experiments.
-Rewarding Doubt improves calibration significantly across all settings
-To evaluate how well our approach improves calibration, we compare it against the zero-shot LLM (Verbalize) and existing methods in both the Single-Answer and Multiple-Answer settings.
-The comparisons for the Single-Answer setting are provided in Table 1, while those for the Multiple-Answer setting are shown in Table 2.
-Across all settings, our approach significantly improves the LLM’s confidence calibration compared to zero-shot verbalization and outperforms most baselines, which often completely fail to produce accurate confidence estimates, in both ECE and AUROC.
-The only exception is the TriviaQA dataset, where the Trained Probe achieves a slightly better ECE score, with both methods achieving near-perfect results.
-However, our approach achieves a higher AUROC score indicating that its confidences are more effective for distinguishing between correct and incorrect answers compared to the Trained Probe.
-Since the Trained Probe is the only other trained method, it is not surprising that it outperforms the non-trained baselines.
-These results further support the claim by Azaria & Mitchell (2023) that the model’s internal state holds information about the truthfulness of statements, which can serve as an indicator of uncertainty.
-However, the non-fine-tuned model appears to be unable to effectively associate this internal state with uncertainty.
-Our approach enables the model to make use of this correlation, resulting in a more accurate expression of the probability that a given answer is correct.
-Notably, our approach consistently outperforms all baselines in the AUROC metrics, even though our reward function only emphasizes calibration accuracy, whereas the Trained Probe is optimized for AUROC.
-As illustrated in Figure 4, the Rewarding Doubt method leads to much improved calibration curves compared to the base model for both TriviaQA and QAMPARI.
-The main outlier on QAMPARI at the confidence level of 0.8 is not representative because of a small sample size in this bin.
-Method ECE ($\downarrow$) AUROC ($\uparrow$)
-Verbalize 0.3459 0.5858
-Chain-of-Thought 0.3065 0.6379
-Surrogate Token 0.3686 0.5923
-Sequence Probability 0.3156 0.7804
-Self-Consistency 0.1134 0.8213
-Trained Probe 0.0189 0.8173
-Rewarding Doubt (ours) 0.0226 0.8592
-Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting.
-(a) Base model
-(b) Fine-tuned by Rewarding Doubt
-Figure 4: Calibration curves of the zero-shot base model (Verbalize) and the model fine-tuned by Rewarding Doubt.
-Method ECE ($\downarrow$) AUROC ($\uparrow$)
-Verbalize 0.5319 0.6047
-Sequence probability 0.3230 0.5942
-Trained probe 0.2119 0.6481
-Rewarding doubt (ours) 0.1061 0.7268
-Table 2: Comparison of methods on the QAMPARI dataset in the Multiple-Answer setting.
-Rewarding Doubt breaks the pattern of overconfidence
-Before fine-tuning the base model, we can observe in Figure 5(a) that the model is very overconfident and almost always reports a confidence of eight or higher.
-This pattern is also observed by Xiong et al.
-(2024), who hypothesize that this overconfidence comes from the supervised pre-training, where the training data usually expresses high confidence, leading the model to mimic these surface-level patterns instead of developing a real sense of confidence.
-After fine-tuning the model with our method, as shown in Figure 5(b), it expresses confidence scores distributed over a much wider range, including lower confidence values.
-This shift indicates that the model is able to overcome the learned pattern of overconfidence and develops an improved sense of confidence.
-Rewarding Doubt shows significant generalization capabilities between datasets
-To assess the generalization capabilities of our method, we evaluate the model trained on TriviaQA in an out-of-domain setting, using the CommonsenseQA and MedQA datasets.
-The results are presented in Table 3 and Table 4, respectively.
-Our approach is significantly better able to generalize than the Trained Probe method, which performs poorly on both datasets and completely fails to generalize to CommonsenseQA.
-Additionally, our fine-tuned model is significantly better calibrated than the base model on MedQA.
-However, on CommonsenseQA, the ECE of our approach is slightly worse than that of the base model.
-This reveals a flaw in the ECE metric.
-Notably, the ECE does not account for the range of confidence values.
-For instance, if a model were to consistently predict a confidence of 6 for a dataset with an overall accuracy of 60%, the ECE would be optimal.
-However, such confidence values would lack practical utility, as they would not enable meaningful discrimination between individual cases.
-Therefore, a low ECE should not be the sole criterion for evaluating confidence estimation.
-A model’s ability to discriminate between correct and incorrect answers, as measured by the AUROC score, should also be considered.
-Consequently, the improvement in the AUROC score compared to the zero-shot verbalization, supports the increased utility of the confidence estimates produced by the fine-tuned model.
-Overall, these generalization experiments demonstrate that the uncertainty indicators the model learns are not entirely dataset-specific; rather, our method enables the model to develop a more generalized understanding of confidence.
-Method ECE ($\downarrow$) AUROC ($\uparrow$)
-Verbalize 0.2820 0.5425
-Trained Probe 0.4819 0.5374
-Rewarding doubt (ours) 0.2930 0.6385
-Table 3: Comparison of generalization results on the CommonsenseQA dataset trained on the TriviaQA dataset.
-Method ECE ($\downarrow$) AUROC ($\uparrow$)
-Verbalize 0.4480 0.5075
-Trained Probe 0.2099 0.5513
-Rewarding doubt (ours) 0.1145 0.6649
-Table 4: Comparison of generalization results on the MedQA dataset trained on the TriviaQA dataset.
-Rewarding Doubt can generalize between settings
-Beyond generalization across datasets, we also assess whether our approach transfers between different experimental settings.
-When evaluating the model trained in the Single-Answer setting directly on a Multiple-Answer task, it has a considerably better calibration compared to the base model, as seen in Table 5.
-While it does not achieve the same level of calibration as the model trained directly in the Multiple-Fact setting, this suggests potential benefits in future applications, particulary for improving the calibration of facts in written texts.
-Training on single facts could already improve the calibration in open-text applications, making it a promising approach for scenarios where structured multi-fact training is impractical.
-Training Evaluation ECE ($\downarrow$) AUROC ($\uparrow$)
-Base model Single fact 0.5875 0.5787
-Single fact Single fact 0.1536 0.7240
-Base model Multi fact 0.5319 0.6047
-Single fact Multi fact 0.1777 0.6617
-Multi fact Multi fact 0.1061 0.7268
-Table 5: Comparison of the base and fine-tuned model on the Qampari dataset in different settings.
-Rewarding Doubt is more resource efficient
-An advantage of our approach, not visible in the metrics, is the inference speed.
-While it initially takes more time for training, inference is highly efficient, as only a small constant number of tokens has to be generated to estimate confidence.
-In constrast, methods like Chain-of-Thought and Self-Consistency need to generate significantly more tokens for their reasoning process, increasing the computational costs and inference time by several orders of magnitude.
-Our approach introduces no additional overhead, does not depend on external resources, and directly provides confidence information through verbalization.
-Thus, the time required to express confidence with our method is significantly less than that of previous approaches.
-(a) Base model
-(b) Fine-tuned by Rewarding Doubt
-Figure 5: Histograms of predicted confidences of the zero-shot base model (Verbalize) and the model fine-tuned on the TriviaQA dataset.
-Rewarding Doubt does not effect answer correctness
-Training confidence calibration with our method only targets the uncertainty estimation abilities and does not effect the performance on the task as seen by the accuracy in Table 6.
-Additionally, our initial hypothesis that the model could learn to generate incorrect answers and label them accordingly with a low confidence score, does not appear to be a problem.
-As shown in Table 6, there is no significant difference in accuracy, calibration, or AUROC scores with and without an additional reward for correct answers.
-Given that our method rewards the model on its expressed confidence and not the answer, it is unlikely that the model deliberately generates incorrect responses.
-Nevertheless, as the inclusion of this additional reward does not negatively impact performance, it can be retained to ensure robust confidence estimation without unintended side effects across different settings.
-Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
-Base model 0.3459 0.5858 0.6310
-Bonus 0.0226 0.8592 0.6309
-No bonus 0.0219 0.8562 0.6278
-Table 6: Comparison of effect of Rewarding Doubt on accuracy on the TriviaQA dataset.
-Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, generalize across tasks, and maintain computational efficiency, making it an effective approach for accurate confidence calibration in language models.
-Limitations
-During our experiments, we observed that the policy occasionally converges to always predicting a fixed confidence value, likely caused by training instabilities.
-While this did not occur consistently, it suggests a potential area for further refinement, such as adjusting the reward structure or introducing additional constraints to encourage a more stable learning of a diverse confidence distribution.
-Additionally, our approach is currently limited to settings, where a binary correctness measure can be established.
-In the future this can be extended to also include continuous correctness measures such as usually applied for free text generation.
-6 Conclusion
-In this work, we propose Rewarding Doubt, a novel approach that enables LLMs to express confidence in their answers more accurately using natural language.
-We leverage reinforcement learning with a reward function that incentivizes diverse and well-calibrated confidence expressions.
-Fine-tuning with our method significantly improves the model’s ability to convey calibrated confidence, effectively reducing the overconfidence commonly observed in LLMs.
-This not only enhances the trustworthiness in AI-generated responses but also lays the groundwork for more reliable human-AI collaboration, where models can transparently communicate uncertainty, an essential step toward safer and more accountable AI systems.
-Acknowledgements
-The authors gratefully acknowledge the financial support by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project ThoraXAI (DIK-2302-0002), and the German Research Foundation (DFG, grant 469106425 - NA 620/51-1).
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Prompts
-For all the question-answering settings, the model is directly prompted to answer a question without a preceding example or context.
-For our method the model was prompted to answer the question and additionally provide a verbalized confidence.
-For the other baselines that do not need a verbalized confidence but infer it indirectly, the model is prompted to only give the correct answer.
-The specifics for multiple-choice are slightly changed but hold mostly the same meaning.
-The exact prompts for each method can be seen in Table 7 for open questions and Table 8 for multiple-choice questions.
-The prompts for each Multi-Answer method can be seen in Table 9.
-We decided not to give the model a role like "expert" in the system prompt but keep it neutral, as we observed that the role we give the model affects the verbalized confidence.
-Method Prompt
-Verbalize/ Rewarding Doubt You will get questions.
-Answer with the correct answer.
-Additionally provide a confidence between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, of how sure you are the answer is correct.
-A value close to 0 means you think there is a high probability that the answer is wrong.
-The closer the value is to 10, the higher you think is the probability that the answer is correct.
-The output should have the format ’Answer: <answer>, Confidence: <confidence>’ and nothing else.
-CoT You will get questions.
-Read the question, analyze step by step, provide your reasoning steps, answer and your confidence in the response.
-The confidence should be between 0, 1, 2, 3, 4, 5.
-A value close to 0 means you think there is a high probability that you could be wrong.
-The closer the value is to 5, the lower you think is the chance that you could be wrong.
-The output should have the format ’Thoughts: <reasoning steps>, Answer: <just the answer and nothing else>, Confidence: <confidence>’.
-Sequence Probability/ Trained Probe You will get questions. Answer with the correct answer only and nothing else.
-[…]
-Self-Consistency You will get questions.
-Read the question, analyze step by step, provide your reasoning steps and answer.
-The output should have the format ’Thoughts: <reasoning steps>, Answer: <just the answer and nothing else>’
-Table 7: The prompts for each method for the open short form questions.
-Method Prompt
-Verbalize/ Rewarding Doubt You will get test questions with possible options.
-Answer with the correct option.
-Additionally provide a confidence between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, of how sure you are the answer is correct.
-A value close to 0 means you think there is a high probability that the answer is wrong.
-The closer the value is to 10, the higher you think is the probability that the answer is correct.
-The output should have the format ’Answer: <answer_index>, Confidence: <confidence>’ and nothing else.
-CoT You will get test questions with possible options.
-Read the question, analyze step by step, provide your reasoningsteps, answer and your confidence in the response.
-The confidence should be between 0, 1, 2, 3, 4, 5.
-A value close to 0 means you think there is a high probability that you could be wrong.
-The closer the value is to 5, the lower you think is the chance that you could be wrong.
-The output should have the format ’Thoughts: <reasoning steps>, Answer: <answer_index>, Confidence: <confidence>’ and nothing else.
-Sequence Probability/ Trained Probe You will get test questions with possible options. Answer with the correct option index only and nothing else.
-[…]
-Self-Consistency You will get test questions with possible options.
-Read the question, analyze step by step, provide your reasoningsteps and the correct option index.
-The output should have the format ’Thoughts: <reasoning steps>, Answer: <answer_index>’ and nothing else.
-Table 8: The prompts for each method for multiple choice questions.
-Method Prompt
-[…]
-[…]
-Table 9: The prompts for each method for multiple fact questions.
+As the logarithm of 0 is undefined, we add a small constant $\epsilon$ in the reward function we use for training:
+$R(a,\hat{p},j)=\begin{cases}log(\max(\hat{p},\epsilon)),&\text{if $j(a)$ = 1 (correct)}\\ log(\min(1-\hat{p},1-\epsilon)),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
+Through this clipping all confidence predictions between 0 and $\epsilon$, and 1 and $1-\epsilon$, respectively, are rewarded equally.
+This leads to the model not being able to differentiate between confidence estimations within these ranges.
+We argue this effect is minor for a sufficiently small $\epsilon$ and can be disregarded in practice.
+Appendix C Societal Impact
+This work introduces a reinforcement learning approach that enables Large Language Models (LLMs) to express calibrated confidence in their factual answers, advancing safe and trustworthy AI deployment.
+The method improves reliability and uncertainty awareness in LLMs, which is particularly valuable in high-stakes settings such as medicine, law, or customer support, where overconfident errors can have serious consequences.
+By optimizing a proper scoring rule during training, our method provides a theoretically sound and generalizable mechanism for aligning confidence with factual correctness—supporting human-AI collaboration and informed decision-making.
+However, expressing numerical confidence may lead users to overly trust AI systems, especially if the model is well-calibrated statistically but still wrong in important individual cases.
+This risk calls for careful deployment, appropriate user interfaces that contextualize model confidence, and safeguards against overreliance on AI-generated outputs.
+Appendix D Asset Details
+The TriviaQA dataset [16] is available at https://nlp.cs.washington.edu/triviaqa/ under the Apache License 2.0.
+The CommonsenseQA dataset [29] (https://www.tau-nlp.sites.tau.ac.il/commonsenseqa) and the MedQA dataset [15] (https://github.com/jind11/MedQA?tab=readme-ov-file) are licensed under the MIT License.
+The QAMPARI dataset [2] is published at https://samsam3232.github.io/qampari/ and licensed under the CC0 1.0 Universal license.
+All these licenses allow broad use, distribution and modification.
+We use the Meta-Llama-3-8B-Instruct [10] language model, provided at https://huggingface.co/unsloth/llama-3-8b-Instruct-bnb-4bit under the Llama 3 Community License Agreement.

### v4

@@
 Empirically, we demonstrate that models trained with our approach exhibit substantially improved calibration and generalize to unseen tasks without further fine-tuning, suggesting the emergence of general confidence awareness.
-We provide our training and evaluation code in the supplementary and will make it publicly available upon acceptance.
+Our code is available at https://github.com/pasta99/RewardingDoubt.
 1 Introduction
-In human intelligence and inter-human interaction, the ability to understand our own uncertainty and communicate our doubts to others is fundamental for effective decision-making, collaboration, and learning [7, 37].
+In human intelligence and inter-human interaction, the ability to understand our own uncertainty and communicate our doubts to others is fundamental for effective decision-making, collaboration, and learning (Cosmides & Tooby, 1996; Xiong et al., 2024).
 Similarly, for Large Language Models (LLMs) to be safely used in real-world applications, especially when humans and AI systems work together, they must not only generate accurate information but also communicate their confidence in that information.
-While LLMs have demonstrated impressive capabilities in natural language understanding, question answering and text summarization [32, 6, 1], LLMs still face significant limitations, such as their tendency to generate inaccurate information, often referred to as hallucinations [11].
+While LLMs have demonstrated impressive capabilities in natural language understanding, question answering and text summarization (Touvron et al., 2023; Chiang et al., 2023; Achiam et al., 2023), LLMs still face significant limitations, such as their tendency to generate inaccurate information, often referred to as hallucinations (Hadi et al., 2023).
 This raises concerns about their reliability, particularly in real-world applications where trustworthiness is essential.
-Especially in high-stakes environments such as medical diagnosis, where LLMs are starting to become support tools for professionals [23, 25, 33, 4], overconfident predictions including factual errors or hallucinations could have serious consequences for patient health.
-Also, in customer service or legal consultation [26, 28], LLMs need to express uncertainty and defer complex queries to human representatives when unsure to avoid misinformed decisions.
+Especially in high-stakes environments such as medical diagnosis, where LLMs are starting to become support tools for professionals (Moor et al., 2023; Pellegrini et al., 2025; Tu et al., 2024; Bani-Harouni et al., 2024), overconfident predictions including factual errors or hallucinations could have serious consequences for patient health.
+Also, in customer service or legal consultation (Shi et al., 2024; Sun et al., 2024), LLMs need to express uncertainty and defer complex queries to human representatives when unsure to avoid misinformed decisions.
 Reliable confidence estimation and expression would enable these systems to flag uncertain outputs for human review, ensuring that crucial decisions are not made based on uncertain LLM outputs.
@@
 Our reward function is designed so that the best reward is given when the confidence and the epistemic probability of being correct are the same, thus incentivizing the expression of calibrated confidences.
-Many previous methods for confidence estimation lack in calibration performance as they do not train the model and instead infer the confidence from the internal state in a zero-shot setup [14, 19, 8].
+Many previous methods for confidence estimation lack in calibration performance as they do not train the model and instead infer the confidence from the internal state in a zero-shot setup (Huang et al., 2023; Kuhn et al., 2023; Duan et al., 2024).
 Additionally, this does not give models an inherent awareness of confidence.
-Other trained methods in this area decouple the uncertainty estimation from the text generation process [3, 18].
+Other trained methods in this area decouple the uncertainty estimation from the text generation process (Azaria & Mitchell, 2023; Kapoor et al., 2024).
 This approach optimizes for calibrated confidence estimation but does not enable the uncertainty-awareness and expression in the model itself.
@@
 We are the first to optimize this function through reinforcement-learning-based policy optimization, leveraging its calibration properties for directly and seamlessly training confidence calibration in LLM generations.
-This reward function captures the fundamental risk-reward tradeoff in probabilistic decision-making, as illustrated in Figure 1.
+This reward function captures the fundamental risk-reward tradeoff in probabilistic decision-making, as illustrated in Figure˜1.
 It increases the reward when a correct answer is given with high confidence, simulating the higher potential return of big bets.
@@
 2 Related Works
-Confidence estimation and calibration have a long history in machine learning and natural language processing [34].
-With the rise of LLMs, a lot of attention has moved to extracting and improving confidence estimation from these new models.
-There is a wide range of methods that can roughly be divided into black-box and white-box methods [9].
-Black-box methods require only the model generations for uncertainty estimation and do not depend on any knowledge of the model’s internal state.
-Within this group, linguistic confidence methods assess the uncertainty by simply prompting the LLM to verbalize its confidence.
-Xiong et al.
-[37] showed that prompting techniques like chain-of-thought (CoT) [36] can, to a limited extend, help the model in the calibration of its confidence expression.
-Based on the assumption that uncertainty in its generation would prime the model to hallucinate information, consistency-based methods [21, 35] measure the similarity between multiple generations and show that high variance in the generation correlates with high uncertainty.
-Wang et al.
-[35] propose using CoT to produce various reasoning pathways towards an answer.
-By marginalizing out the reasoning pathways, they provide the most consistent answer as the one most probably correct and offer the consistency measure as an estimation of the model’s confidence.
+2.1 Confidence Estimation in LLMs
+Confidence estimation and calibration have a long history in machine learning and natural language processing (Wang, 2024).
+With the rise of LLMs, research has focused on adapting and extending these ideas to modern architectures.
+Broadly, methods fall into black-box and white-box approaches (Geng et al., 2024).
+Black-box methods
+Black-box methods rely only on model outputs.
+Linguistic prompting methods ask the model to verbalize its confidence, sometimes aided by chain-of-thought reasoning (Xiong et al., 2024; Wei et al., 2022).
+Consistency-based approaches estimate confidence by measuring agreement across multiple generations, with high variance indicating uncertainty (Manakul et al., 2023; Wang et al., 2022).
+Recently, Zhou et al.
+(2025) proposed SteerConf, which does multiple inference passes where the LLM is prompted to use different levels of caution in its confidence expression.
+The resulting verbalized confidences are aggregated based on confidence and answer consistency to an overall confidence prediction.
 Black-box methods are valuable for their simplicity, ease-of-use and universality, however generally lack behind white-box methods in their calibration performance.
-White-box methods, on the other hand, depend on access to the internal state of the model, either for the evaluation of logits and activations or for fine-tuning of the model for calibration.
-Logit-based methods calibrate the confidence by assessing the probabilities or the entropy of tokens in the predicted sequence.
-Huang et al.
-[14] calculate the mean probability of the generated tokens with the intuition that a model that generates an answer with high probability tokens also has a high confidence in this answer.
-Kuhn et al.
-[19] and [8] expand on this idea by looking at the entropy of the generated tokens while also taking the semantic relevance into account.
-Kadavath et al.
-[17] showed that LLMs can self-evaluate the correctness of their own generation.
-They prompt the model to provide an answer followed by a judgement whether its own answer is "true" or "false".
+White-box methods
+White-box methods exploit internal model states.
+Logit-based techniques estimate confidence from token probabilities or entropy (Huang et al., 2023; Kuhn et al., 2023; Duan et al., 2024), assuming that high probability tokens correspond to high confidence predictions.
+Self-evaluation methods let the model judge the truth of its own answers (Kadavath et al., 2022).
+They prompt the model to provide an answer followed by a judgment whether its own answer is "true" or "false".
 They then compare the probability of the "true" or "false" token to calculate a confidence estimation.
-While all the aforementioned methods are zero-shot methods, another line of white-box methods, proposes to train external models for confidence estimation.
-Xiong et al.
-[37] propose a method that trains a linear probe for assessing the model’s confidence based on the activations of the LLM’s hidden state.
-They evaluate the application of the probe to various layers and train it with correctness labels of the task.
-This approach does not explicitly train on the granularity of confidence estimation and, therefore, is only indirectly able to express a confidence gradient for fine-grained calibration.
-Building on this approach, Mielke et al.
-[22] give an LLM access to the confidence extracted by the external probe and fine-tune it to express this confidence.
-However, the calibration performance is upper-bounded by the ability of the trained probe to assess the model’s confidence accurately.
+External probing approaches train classifiers on hidden states to predict correctness (Azaria & Mitchell, 2023).
+While some of these methods achieve good confidence estimation results, they do not teach the model to express clear confidence values itself but depend on some auxiliary estimation mechanism.
+2.2 Finetuned Confidence Expression
+A growing line of work integrates confidence estimation into instruction tuning.
+These methods typically follow a two-step paradigm: First, they estimate model confidence using various methods, e.g., self-consistency (Cheng et al., 2024; Yang et al., 2024; Han et al., 2024), token probabilities (Chen et al., 2024), trained probes (Mielke et al., 2022), empirical accuracy (Zhang et al., 2024; Lin et al., 2022; Ulmer et al., 2024), or topic unfamiliarity (Wan et al., 2024; Kang et al., 2024).
+Second, they construct finetuning datasets that either replace uncertain answers with refusals (Zhang et al., 2024; Cheng et al., 2024; Yang et al., 2024; Wan et al., 2024) or append the estimated uncertainty as an additional supervised signal (Han et al., 2024; Chen et al., 2024; Mielke et al., 2022; Lin et al., 2022; Ulmer et al., 2024).
+The key limitation of this approach is that the model’s expressed confidence is bounded by the quality of the constructed ground-truth estimates.
+Additionally, while the underlying confidence estimation method might optimize for perfect calibration (e.g.
+in the case of the trained probe), this theoretical guarantee is lost when performing supervised finetuning on these constructed ground truths to reproduce these scores.
+2.3 Reinforcement Learning for Confidence Expression
+Reinforcement Learning from Human Feedback (RLHF) has proven effective for aligning LLMs with human preferences (Ouyang et al., 2022), and has also been explored for agentic interaction in textual environments (Zhou et al., 2023; Carta et al., 2023).
+Only recently have researchers begun applying RL directly to confidence estimation.
+Tao et al.
+(2024) adapt RLHF by designing rewards that align verbalized confidence with preference ratings, but this requires human-annotated preference data and does not address factual calibration.
+Leng et al.
+(2024) identify that standard reward models in RLHF are biased toward high verbalized confidence, rating answers with high confidence expressions with a high reward.
+To counteract this, they introduce two reward model training paradigms, PPO-M and PPO-C, which fine-tune the reward model to reward answers where correctness and confidence expression are aligned.
+Xu et al.
+(2024) propose RL from Knowledge Feedback (RLKF) to encourage refusals outside the model’s knowledge scope, reducing hallucinations but without quantifying confidence.
 Stengel-Eskin et al.
-[27] propose a DPO-based method, LACIE, simulating an interaction between a generation and a listener model, where the speaker model generates answers, and the listener model evaluates them based on perceived confidence.
-By training with Direct Preference Optimization (DPO) on data from TriviaQA, LACIE rewards accurate confidence expression, including correct answers accepted by the listener model and incorrect, rejected ones, while penalizing misleading confidences.
-The confidence expression is calibrated to how the listener model interprets confidence cues, depending on its ability to judge confidence and aligning perceived confidence rather than a fact-based numerical confidence score.
-However, none of these methods provide theoretical guarantees or provable optimality regarding the trained calibration of the expressed confidence.
-Simultaneously, fine-tuning LLMs using Reinforcement Learning from human feedback (RLHF) has received increasing attention for aligning the models with human preferences, avoiding toxicity or the generation of harmful content [24, 20].
-Outside of RLHF, RL has been used to fine-tune LLMs for agentic interaction in textual environments [40, 5].
-Confidence estimation of LLMs using RL is a novel research direction that, to the best of our knowledge, has only been approached by Tao et al.
-[30] and Xu et al.
-[38].
-Tao et al.
-[30] build on the Reinforcement Learning with Human Feedback (RLHF) paradigm and design a reward function that aligns the verbalized confidence of the model with the preference rating predicted by the RLHF reward model.
-However, this approach can not evaluate model confidence in factual statements and depends on human-annotated preference data.
-Xu et al.
-[38] introduce a novel Reinforcement Learning from Knowledge Feedback (RLKF) framework, which aims to teach a model to refuse to answer questions outside their knowledge scope, reducing hallucinations, but does not quantify confidence when giving an answer.
-In contrast, our method directly optimizes for factual calibration using a theoretically grounded, proper scoring rule as the reward signal, enabling the model to develop intrinsic uncertainty awareness without requiring external preference models, knowledge supervision, or post-hoc calibration techniques, while at the same time seamlessly integrating calibrated confidence expression into the LLMs response generation.
+(2024) propose LACIE, a DPO-based approach that simulates an interaction between a speaker and a listener model, rewarding accurate and honest confidence expression by aligning it with the listener’s interpretation of confidence cues rather than with fact-based numerical calibration.
+In contrast to previous works, our method directly optimizes for factual calibration using a theoretically grounded, proper scoring rule as the reward signal, enabling the model to develop intrinsic uncertainty awareness without requiring external preference models, knowledge supervision, or post-hoc calibration techniques, while at the same time seamlessly integrating calibrated confidence expression into the LLMs response generation.
 3 Rewarding Doubt
 We propose a novel reinforcement-learning approach, that improves an LLM’s ability to verbalize an accurate numerical confidence in a previously generated answer.
-The LLM functions as an agent in a simulated environment as shown in Figure 2, that poses challenging question-answering scenarios.
+The LLM functions as an agent in a simulated environment as shown in Figure˜2, that poses challenging question-answering scenarios.
 It is prompted with task queries such as factual questions and asked to predict both an answer to the query as well as a confidence score.
@@
 Correct answers with high confidences are rewarded highly, but also penalized heavily when incorrect.
-Formally, let the model be provided with a textual question or request $q$, resulting in an answer-confidence pair $(a,\hat{p})$ as response, where $a$ is a textual answer with binary correctness value, and $0\leq\hat{p}\leq 1$ is a numerical confidence score representing the subjective probability the model assigns to answer $a$ being correct.
-We train this subjective probability assessment to align with the true epistemic probability $p^{*}$, which represents the actual likelihood of correctness given the model’s internal knowledge state.
-If $\hat{p}$ and $p^{*}$ are aligned the model is perfectly calibrated, meaning the probability of correctness $P(j(a)=1)$ always equals the expressed confidence:
+[…]
 $P(j(a)=1\mid\hat{p}=x)=x,\quad\forall x\in[0,1],$
@@
   • Reward function ($R$): The reward $R(a,c,j)$ is computed based on the final confidence score sequence $c=(c_{1},\dots,c_{T})$ and the correctness of the answer $j(a)$.
-To promote accurate confidence estimation, the model’s reward function must fulfill the requirement of being maximized when $\hat{p}=p^{*}$.
-The model should receive a higher reward when it correctly predicts an outcome with high confidence.
-Conversely, incorrect predictions should be penalized more severely when the model exhibits high confidence.
-In contrast, expressing uncertainty reduces potential penalties, leading to a higher expected reward in uncertain scenarios.
+To promote accurate confidence estimation, the model’s expected reward must fulfill the requirement of being maximized when $\hat{p}=p^{*}$, i.e.
+when the predicted confidence aligns with the probability of correctness.
+The model should receive a high reward when it correctly predicts an outcome with high confidence or when it incorrectly predicts an outcome with low confidence.
+Conversely, the reward should be low when incorrect predictions are provided with high confidence or correct predictions are provided with low confidence.
 This approach incentivizes the model to express high confidence only in cases where certainty is warranted while expressing doubt in ambiguous situations.
 By penalizing both overconfidence and underconfidence, the model is encouraged to calibrate its confidence accurately, effectively balancing the trade-off between reward maximization and penalty avoidance.
-We design our reward as a logarithmic scoring function, fulfilling the described optimality conditions:
-$R(a,\hat{p},j)=\begin{cases}log(\max(\hat{p},\epsilon)),&\text{if $j(a)$ = 1 (correct)}\\ log(\min(1-\hat{p},1-\epsilon)),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
-where $\epsilon>0$ is a small positive constant introduced for numerical stability to avoid evaluating the logarithm at zero.
-The normalized reward for correct and incorrect answers for each confidence is visualized in Figure 3.
+Note, that through this design our method focuses exclusively on improving calibration while keeping task performance stable.
+We design our reward as a logarithmic scoring function:
+[…]
+This function fulfills the requirement described above as we show in the following proposition:
 Proposition 1 (Optimality implies Calibration).
@@
 The proof of Proposition 1 is analogous to the proof that the logarithmic scoring rule is a proper scoring rule.
-We provide it in full in Appendix B and discuss the influence of the clipping on the optimality of the reward function.
+We provide it in full in Appendix˜B and discuss the influence of the clipping on the optimality of the reward function.
+Since the logarithm of zero is undefined, we introduce a small positive constant $\epsilon$ as clipping value for numerical stability.
+Concretely, we clip the lower and upper limit of the confidence $\hat{p}$ to $\epsilon$ and $1-\epsilon$, respectively.
+The clipped reward function is provided in Appendix˜C.
+The normalized and clipped reward for correct and incorrect answers for each confidence is visualized in Figure˜3.
 Figure 3: The rewards for each confidence value for correct and incorrect answers.
@@
 A confidence of zero is defined as the model being certain that the answer is incorrect, while ten is defined as the model being certain the answer is correct.
-We normalize the reward function to the range of $[-1,1]$.
-In the Single-Answer setting we train the model on the TriviaQA dataset [16], which contains question-answer-evidence triplets, from which we only use the questions and answers.
-For generalization experiments, we evaluate our method on CommonsenseQA [29] and MedQA [15], which are multiple-choice question datasets in the commonsense and medical domain, respectively.
-For the Multiple-Answer setting, we train on the QAMPARI dataset [2], which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
-In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought [37], Top-K [31], Surrogate Token [17], Sequence Probability [14] and Self-Consistency [35] as zero-shot methods, LACIE [27], which uses DPO for optimizing confidence expression and Trained Probe [3], which employs supervised training of an external probe for estimation model.
+We normalize the reward function to the range of [1,1][-1,1].
+In the Single-Answer setting we train the model on the TriviaQA dataset (Joshi et al., 2017), which contains question-answer-evidence triplets, from which we only use the questions and answers.
+For generalization experiments, we evaluate our method on CommonsenseQA (Talmor et al., 2019) and MedQA (Jin et al., 2020), which are multiple-choice question datasets in the commonsense and medical domain, respectively.
+For the Multiple-Answer setting, we train on the QAMPARI dataset (Amouyal et al., 2023), which contains questions with multiple-answers as well as evidence, again only using the questions and answers.
+In the Single-Answer setting we compare our approach on the TriviaQA dataset against the following methods: Chain-of-Thought (Xiong et al., 2024), Top-K (Tian et al., 2023), Surrogate Token (Kadavath et al., 2022), Sequence Probability (Huang et al., 2023) and Self-Consistency (Wang et al., 2022) as zero-shot methods, LACIE (Stengel-Eskin et al., 2024), which uses DPO for optimizing confidence expression and Trained Probe (Azaria & Mitchell, 2023), which employs supervised training of an external probe for estimation model.
 We also compare to the non-finetuned base model in a zero-shot manner, using the same prompt as our Rewarding Doubt method and refer to this setup as Verbalize.
@@
 We report our results using the Expected Calibration Error (ECE) and the Area Under the Receiver Operating Characteristic Curve (AUROC) metric.
-The ECE measures the difference between a model’s predicted confidence and its actual accuracy, assessing how well confidences align with real outcomes.
-AUROC assesses a model’s ability to distinguish between classes, with a score of 0.5 indicating random guessing and 1.0 indicating perfect classification.
 Additionally, we visualize the calibration with calibration curves, where a well-calibrated model lies close to the 45° line and large deviations show a high miscalibration.
@@
 We consider an answer as correct if its score exceeds a threshold of 0.5.
-5 Implementation Details
-Training
-As base model for our experiments we use Meta-Llama-3-8B-Instruct [10].
-We employ the 4-bit quantized performance-optimized model version by Unsloth AI [12].
-Additionally, we apply LoRA fine-tuning [13], reducing the amount of trainable parameters from 8,030,265,345 to 41,947,137.
+Implementation Details
+We optimize the reward function using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017).
+Unless stated otherwise, we use Meta-Llama-3-8B-Instruct (Grattafiori et al., 2024) as base model for our experiments.
+We employ the 4-bit quantized performance-optimized model version by Unsloth AI (Han et al., 2023) and apply LoRA fine-tuning (Hu et al., 2022).
 For the Single-Answer setting we train the model for two epochs with a learning rate of 1e-5.
-For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 24,000 steps with a batchsize of eight and a learning rate of 1e-5 with a reward scale of five.
+For the Multiple-Answer setting, due to the size of the training dataset and the fact that each question yields multiple facts, the model is trained for a limited amount of 24,000 steps with a batchsize of eight and a learning rate of 1e-5 and multiply the reward with 5 to increase its spread.
 All models are trained on one Nvidia A40 with each training run taking seven days.
@@
 If the model fails to generate an answer in the specified format, it is penalized with an out-of-format reward of -3.
-Baselines
-For the Sequence Probability, we compute the average probability for each token in the response.
-In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric [39].
-For the trained probe [3], the original study introduced a custom dataset comprising short statements classified as either true or false.
-The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements.
-To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers.
-The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework.
-For the architecture of the MLP, we employed the same design as Azaria and Mitchell [3] and train it for four epochs with a learning rate of 1e-4 until convergence.
-The exact prompts used for each baseline are provided in Appendix A.
+Detailed implementation choices for the baselines are provided in Appendix˜D.
 5 Results and Discussion
-Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting. *Results are from the original paper [27].
+This section presents and discusses the key findings of our experiments for both Single and Multiple-Answer tasks and the generalization to out-of-domain datasets.
+Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting with 95% CIs in brackets.
+* Results are from the original paper (Stengel-Eskin et al., 2024) and include standard error.
 Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
-Verbalize 0.3459 0.5858 0.6310
-Chain-of-Thought 0.3065 0.6379 0.6273
-Top-K 0.1611 0.6673 0.6023
-Surrogate Token 0.3686 0.5923 0.5933
-Sequence Probability 0.3156 0.7804 0.5955
-Self-Consistency 0.1134 0.8213 0.6224
-LACIE* 0.1200 0.7200 n/a
-Trained Probe 0.0189 0.8173 0.5925
-Rewarding Doubt (ours) 0.0226 0.8592 0.6309
-This section presents and discusses the key findings of our experiments for both Single and Multiple-Answer tasks and the generalization to out-of-domain datasets.
-Calibration Performance
+Verbalize 0.3459 [0.3375,0.3543] 0.5858 [0.5778,0.5936] 0.6310 [0.6222,0.6397]
+Chain-of-Thought 0.3065 [0.2981,0.3157] 0.6379 [0.6284,0.6475] 0.6273 [0.6181,0.6363]
+Top-K 0.1611 [0.1529,0.1695] 0.6673 [0.6580,0.6768] 0.6023 [0.5936,0.6110]
+Surrogate Token 0.3686 [0.3595,0.3783] 0.5923 [0.5818,0.6027] 0.5933 [0.5844,0.6016]
+Sequence Probability 0.3156 [0.3074,0.3237] 0.7804 [0.7725,0.7876] 0.5955 [0.5864,0.6040]
+Self-Consistency 0.1134 [0.1066,0.1210] 0.8213 [0.8129,0.8298] 0.6224 [0.6131,0.6317]
+PPO-M 0.3262 [0.3173,0.3346] 0.5274 [0.5227,0.5319] 0.5749 [0.5662,0.5835]
+PPO-C 0.3607 [0.3524,0.3697] 0.5439 [0.5384,0.5491] 0.5258 [0.5164,0.5358]
+LACIE* 0.1200 $\pm$0.02 0.7200 $\pm$0.02 n/a
+Trained Probe 0.0189 [0.0147,0.0275] 0.8173 [0.8099,0.8250] 0.5925 [0.5834,0.6017]
+Rewarding Doubt (ours) 0.0226 [0.0176,0.0302] 0.8592 [0.8523,0.8664] 0.6309 [0.6222,0.6399]
 To assess how well our approach improves calibration, we compare it against the zero-shot LLM baseline (Verbalize) and several established methods in both Single-Answer and Multiple-Answer question-answering tasks.
-Results for the Single-Answer setting on TriviaQA are presented in Table 1, and those for the Multiple-Answer setting on QAMPARI appear in Table 2.
+Results for the Single-Answer setting on TriviaQA are presented in Table˜1, and those for the Multiple-Answer setting on QAMPARI appear in Table˜2.
 Across both tasks, Rewarding Doubt substantially improves the model’s confidence calibration over zero-shot verbalization.
@@
 In comparison, Verbalize, Sequence Probability, and Trained Probe perform notably worse.
-Our findings support the claim by Azaria and Mitchell [3] that a model’s internal state encodes information about the truthfulness of statements, which can serve as an indicator of uncertainty.
+Our findings support the claim by Azaria & Mitchell (2023) that a model’s internal state encodes information about the truthfulness of statements, which can serve as an indicator of uncertainty.
 However, without fine-tuning, the model struggles to utilize this internal information effectively.
 Our approach enables the model to make use of this correlation and translate it into an accurate expression of the probability that a given answer is correct.
-The calibration curves in Figure 4 further illustrate these improvements.
+The calibration curves in Figure˜4 further illustrate these improvements.
 For both TriviaQA and QAMPARI, the fine-tuned model’s confidence much more closely aligns with the ideal 45° line than the zero-shot Verbalize baseline.
 Additionally, we observe a shift in the confidence distribution after fine-tuning.
-As shown in Figure 5, in a zero-shot setting the LLM (Verbalize) predominantly assigns high confidence scores (8 or above), reflecting overconfidence, a pattern also noted by Xiong et al.
-[37], who attribute it to supervised pretraining that favors confident expressions.
-After fine-tuning with Rewarding Doubt, the model’s confidence scores (shown in Figure 5(b)) span a wider range, including lower values, indicating a more nuanced expression of uncertainty.
+As shown in Figure˜5, in a zero-shot setting the LLM (Verbalize) predominantly assigns high confidence scores (8 or above), reflecting overconfidence, a pattern also noted by Xiong et al.
+(2024), who attribute it to supervised pretraining that favors confident expressions.
+After fine-tuning with Rewarding Doubt, the model’s confidence scores (shown in Figure˜5(b)) span a wider range, including lower values, indicating a more nuanced expression of uncertainty.
 This shift suggests that fine-tuning mitigates overconfidence and better aligns the model’s confidence with its actual performance.
@@
 Figure 4: Calibration curves of the zero-shot base model (Verbalize) and the model fine-tuned by Rewarding Doubt.
-Table 2: Comparison of methods on the QAMPARI dataset in the Multiple-Answer setting.
+Table 2: Comparison of methods on the QAMPARI dataset in the Multiple-Answer setting with 95% CIs in brackets.
 Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
-Verbalize 0.5319 0.6047 0.2549
-Sequence probability 0.5324 0.5942 0.1927
-Trained probe 0.1117 0.6481 0.2233
-Rewarding doubt (ours) 0.0816 0.6947 0.2480
+Verbalize 0.5319 [0.5172,0.5461] 0.6047 [0.5837,0.6267] 0.2550 [0.2410,0.2698]
+Sequence probability 0.5324 [0.5225,0.5432] 0.5942 [0.5775,0.6110] 0.1928 [0.1829,0.2024]
+Trained probe 0.1117 [0.0997,0.1262] 0.6481 [0.6241,0.6726] 0.2233 [0.2094,0.2384]
+Rewarding doubt (ours) 0.0816 [0.0723,0.0951] 0.6947 [0.6776,0.7113] 0.2480 [0.2348,0.2609]
 (a) Base model
@@
 Figure 5: Histograms of predicted confidences of the zero-shot base model (Verbalize) and the model fine-tuned on the TriviaQA dataset.
-Table 3: Comparison of generalization results on the CommonsenseQA dataset trained on the TriviaQA dataset.
-Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
-Verbalize 0.2820 0.5425 0.6860
-Trained Probe 0.4819 0.5374 0.7108
-Rewarding doubt (ours) 0.2930 0.6385 0.7163
+To test the consistency of our method across different models, we perform an ablation study across diverse LLM architectures and sizes.
+Specifically, we apply Rewarding Doubt to Qwen-2.5 (3B and 7B) and Gemma-2 (9B) models, in addition to LLaMA-3.1-8B.
+Table 3 reports performance for each model before and after fine-tuning with our method.
+Despite architectural and pretraining differences, Rewarding Doubt consistently reduces calibration error and improves AUROC across all models, without degrading downstream accuracy.
+Table 3: Calibration and accuracy of Verbalize vs. Rewarding Doubt across different LLMs with 95% CIs in brackets.
+Model Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
+LLaMA-3.1-8B Verbalize 0.2771 [0.2689,0.2862] 0.6766 [0.6667,0.6863] 0.6662 [0.6577,0.6745]
+Trained probe 0.0152 [0.0118,0.0235] 0.8495 [0.8420,0.8567] 0.6231 [0.6143,0.6322]
+Rew. Doubt 0.0256 [0.0209,0.0327] 0.8793 [0.8729,0.8860] 0.6497 [0.6407,0.6585]
+Qwen-2.5-3B Verbalize 0.5330 [0.5252,0.5435] 0.5981 [0.5927,0.6035] 0.4185 [0.4085,0.4247]
+Trained probe 0.0186 [0.0134,0.0268] 0.7975 [0.7880,0.8066] 0.2540 [0.2463,0.2624]
+Rew. Doubt 0.1483 [0.1415,0.1546] 0.9065 [0.9012,0.9122] 0.4193 [0.4097,0.4283]
+Qwen-2.5-7B Verbalize 0.3619 [0.3530,0.3705] 0.5818 [0.5762,0.5879] 0.5239 [0.5148,0.5331]
+Trained probe 0.0989 [0.0920,0.1057] 0.8737 [0.8676,0.8797] 0.4793 [0.4696,0.4881]
+Rew. Doubt 0.1298 [0.1237,0.1368] 0.8928 [0.8866,0.8988] 0.5283 [0.5193,0.5368]
+Gemma-2-9B Verbalize 0.3206 [0.3122,0.3288] 0.5615 [0.5548,0.5682] 0.6690 [0.6603,0.6773]
+Trained probe 0.0301 [0.0253,0.0373] 0.8694 [0.8629,0.8769] 0.6464 [0.6380,0.6551]
+Rew. Doubt 0.0922 [0.0861,0.0994] 0.8649 [0.8570,0.8725] 0.6832 [0.6743,0.6918]
 Stability of Answer Correctness
@@
 Generalization Capabilities
-To assess the generalization abilities of Rewarding Doubt, we evaluated the model trained on TriviaQA in out-of-domain settings using the CommonsenseQA [29] and MedQA [15] datasets.
-Results are shown in Table 3 and Table 4, respectively.
+To assess the generalization abilities of Rewarding Doubt, we evaluated the model trained on TriviaQA in out-of-domain settings using the CommonsenseQA (Talmor et al., 2019) and MedQA (Jin et al., 2020) datasets.
+Results are shown in Table˜4.
 On MedQA, Rewarding Doubt significantly outperforms Verbalize in both metrics, while on CommonsenseQA, it achieves a comparable ECE, however paired with a much higher AUROC.
@@
 A model consistently assigning moderate confidence values could appear well-calibrated under ECE, yet fail to offer meaningful distinctions between uncertain and certain cases.
-The AUROC, by contrast, directly measures this discriminative ability.
+AUROC, in contrast, directly measures this discriminative ability.
 Thus, the substantial improvements in AUROC underscore that Rewarding Doubt produces more useful and actionable confidence estimates.
-Compared to the Trained Probe baseline, Rewarding Doubt consistently outperforms with significantly improved metrics, showing a stronger ability to generalize to new datasets.
-Table 4: Comparison of generalization results on the MedQA dataset trained on the TriviaQA dataset.
+Compared to the Trained Probe, the best-performing baseline, Rewarding Doubt consistently outperforms, showing a stronger ability to generalize to new datasets.
+Table 4: Comparison of generalization results on CommonsenseQA (CsQA) and MedQA, trained on the TriviaQA dataset with 95% CIs in brackets.
 Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
-Verbalize 0.4480 0.5075 0.5066
-Trained Probe 0.2099 0.5513 0.5051
-Rewarding doubt (ours) 0.1145 0.6649 0.5161
-We also explore generalization across experimental settings by applying a model trained in a Single-Answer setting to a Multiple-Answer task.
-Although it does not reach the performance of a model trained specifically for that task, it still outperforms baseline methods considerably, demonstrating transferability of the learning confidence estimation patterns.
+CsQA Verbalize 0.2820 [0.2206,0.3422] 0.5425 [0.4740,0.6069] 0.6860 [0.6277,0.7444]
+Trained Probe 0.4819 [0.4655,0.5130] 0.5374 [0.5021,0.5708] 0.7108 [0.6847,0.7355]
+Rewarding doubt (ours) 0.2930 [0.2693,0.3179] 0.6385 [0.6065,0.6715] 0.7163 [0.6918,0.7410]
+MedQA Verbalize 0.4480 [0.4200,0.4753] 0.5075 [0.4803,0.5338] 0.5067 [0.4784,0.5350]
+Trained Probe 0.2099 [0.1881,0.2439] 0.5513 [0.5207,0.5844] 0.5051 [0.4792,0.5318]
+Rewarding doubt (ours) 0.1145 [0.0893,0.1408] 0.6649 [0.6355,0.6954] 0.5161 [0.4886,0.5420]
+We also explore generalization across experimental settings in Table˜5 by applying a model trained in a Single-Answer setting to a Multiple-Answer task.
+Although under-performing a model trained specifically for that task, it still outperforms the base model considerably, demonstrating transferability of the learned confidence estimation patterns.
 This suggests promising applications for improving confidence estimation in more complex or less structured scenarios, such as fact verification and calibration in free-text generation, even when specialized training data is unavailable.
-Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, while generalizing across tasks, and maintaining stable task performance, making it an effective approach for accurate confidence calibration and expression in language models.
-Beyond improvements in calibration quality, our method also offers practical advantages.
+Our current experiments focus on settings where answer quality can be evaluated via exact rule-based metrics, yielding a binary correctness signal, the Rewarding Doubt framework could be extended to work with correctness signals provided by an LLM-as-a-judge system, a reward model trained on human preferences or continuous NLG metrics.
+Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, while generalizing across tasks, and maintaining stable task performance, making it an effective approach for accurate confidence calibration and expression in LLMs.
+Beyond improvements in calibration, our method also offers practical advantages.
 While fine-tuning requires an initial training investment, inference remains highly efficient, as only a small, constant number of tokens need to be generated to express confidence.
 In contrast, zero-shot methods like Chain-of-Thought and Self-Consistency have substantial computational overhead during inference by requiring lengthy reasoning chains or multiple generations.
-Rewarding Doubt introduces no such overhead, does not rely on an additional model, and directly provides actionable confidence estimates through simple verbalization directly by the LLM, making it highly suitable for real-world deployment.
-Table 5: Comparison of the base and fine-tuned model on the Qampari dataset in different settings.
+Rewarding Doubt introduces no such overhead, does not rely on an additional model, and directly provides actionable confidence estimates through verbalization directly by the LLM, making it highly suitable for real-world deployment.
+Table 5: Comparison of the base and fine-tuned model on the Qampari dataset in different settings with 95% CIs in brackets.
 Training Evaluation ECE ($\downarrow$) AUROC ($\uparrow$)
-Base model Single fact 0.5875 0.5787
-Single fact Single fact 0.1536 0.7240
-Base model Multi fact 0.5319 0.6047
-Single fact Multi fact 0.1777 0.6617
-Multi fact Multi fact 0.1061 0.7268
+Base model Single fact 0.5875 [0.5597,0.6151] 0.5787 [0.5408,0.6125]
+Single fact Single fact 0.1536 [0.1320,0.1813] 0.7240 [0.6889,0.7577]
+Base model Multi fact 0.5319 [0.5172,0.5461] 0.6047 [0.5837,0.6267]
+Single fact Multi fact 0.1777 [0.1679,0.1890] 0.6617 [0.6468,0.6779]
+Multi fact Multi fact 0.1061 [0.0935,0.1206] 0.7268 [0.7065,0.7468]
 Limitations
-During our experiments, we observed that the policy occasionally converges to always predicting a fixed confidence value, likely caused by training instabilities.
-While this did not occur consistently, it suggests a potential area for further refinement, such as adjusting the reward structure or introducing additional constraints to encourage a more stable learning of a diverse confidence distribution.
-Additionally, our approach is currently limited to settings, where a binary correctness measure can be established.
-In the future this can be extended to also include continuous correctness measures such as usually applied for free text generation.
+Due to computational constraints, we only tested Rewarding Doubt on models with sizes ranging from 3B to 9B parameters.
+While we expect similar effectiveness on larger models, empirical validation on such scales would be valuable.
 6 Conclusion
@@
 The authors gratefully acknowledge the financial support by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project ThoraXAI (DIK-2302-0002), and the German Research Foundation (DFG, grant 469106425 - NA 620/51-1).
+Reproducibility Statement
+In order to ensure reproducibility, we describe implementation details of Rewarding Doubt as well as the used baselines in Section˜4 and Appendix˜D.
+Further, Appendix˜A provides the exact prompts used for different experiments.
+Lastly, we publish our code at https://github.com/pasta99/RewardingDoubt.
+Appendix
 Appendix A Prompts
@@
 The specifics for multiple-choice are slightly changed but hold mostly the same meaning.
-The exact prompts for each method can be seen in Table 6 for open questions and Table 7 for multiple-choice questions.
-The prompts for each Multi-Answer method can be seen in Table 8.
+The exact prompts for each method can be seen in Table˜6 for open questions and Table˜7 for multiple-choice questions.
+The prompts for each Multi-Answer method can be seen in Table˜8.
 We decided not to give the model a role like "expert" in the system prompt but keep it neutral, as we observed that the role we give the model affects the verbalized confidence.
@@
 In the following, we prove Proposition 1 with the reward function
-$R(a,\hat{p},j)=\begin{cases}log(\hat{p}),&\text{if $j(a)$ = 1 (correct)}\\ log(1-\hat{p}),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
+[…]
 Proof.
@@
 Let $f(\hat{p})=\mathbb{E}[R(a,\hat{p},j)]$ be the expected reward for all values of $\hat{p}$ and $p^{*}$:
-$f(\hat{p})=p^{*}\log(\hat{p})\;+\;(1-p^{*})\log\bigl{(}1-\hat{p}\bigr{)}.$
+[…]
 Taking the first derivative w.r.t. $\hat{p}$:
@@
 $f^{\prime\prime}(\hat{p})=-\frac{p^{*}}{\hat{p}^{2}}\;-\;\frac{1-p^{*}}{\bigl(1-\hat{p}\bigr)^{2}}$
-is strictly negative for $\hat{p}\in(0,1)$. Hence, $f(\hat{p})$ is concave and has its global maximum at $\hat{p}=p^{*}$. ∎
+is strictly negative for $\hat{p}\in(0,1)$. Hence, f(p^)f($\hat${p}) is concave and has its global maximum at $\hat{p}=p^{*}$. ∎
 As the logarithm of 0 is undefined, we add a small constant $\epsilon$ in the reward function we use for training:
-$R(a,\hat{p},j)=\begin{cases}log(\max(\hat{p},\epsilon)),&\text{if $j(a)$ = 1 (correct)}\\ log(\min(1-\hat{p},1-\epsilon)),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
+[…]
 Through this clipping all confidence predictions between 0 and $\epsilon$, and 1 and $1-\epsilon$, respectively, are rewarded equally.
@@
 We argue this effect is minor for a sufficiently small $\epsilon$ and can be disregarded in practice.
-Appendix C Societal Impact
+Appendix C Clipped Reward Function
+The clipped reward function as described in Section˜3, is defined as follows:
+[…]
+where $\epsilon>0$ is a small positive constant of 0.001 introduced for numerical stability to avoid evaluating the logarithm at zero.
+Appendix D Implementation Details of Baselines
+For the Sequence Probability, we compute the average probability for each token in the response.
+In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric Zhang et al.
+(2019).
+For the trained probe Azaria & Mitchell (2023), the original study introduced a custom dataset comprising short statements classified as either true or false.
+The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements.
+To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers.
+The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework.
+For the architecture of the MLP, we employed the same design as Azaria & Mitchell (2023) and train it for four epochs with a learning rate of 1e-4 until convergence.
+The exact prompts used for each baseline are provided in Appendix˜A.
+Appendix E Societal Impact
 This work introduces a reinforcement learning approach that enables Large Language Models (LLMs) to express calibrated confidence in their factual answers, advancing safe and trustworthy AI deployment.
@@
 This risk calls for careful deployment, appropriate user interfaces that contextualize model confidence, and safeguards against overreliance on AI-generated outputs.
-Appendix D Asset Details
-The TriviaQA dataset [16] is available at https://nlp.cs.washington.edu/triviaqa/ under the Apache License 2.0.
-The CommonsenseQA dataset [29] (https://www.tau-nlp.sites.tau.ac.il/commonsenseqa) and the MedQA dataset [15] (https://github.com/jind11/MedQA?tab=readme-ov-file) are licensed under the MIT License.
-The QAMPARI dataset [2] is published at https://samsam3232.github.io/qampari/ and licensed under the CC0 1.0 Universal license.
-All these licenses allow broad use, distribution and modification.
-We use the Meta-Llama-3-8B-Instruct [10] language model, provided at https://huggingface.co/unsloth/llama-3-8b-Instruct-bnb-4bit under the Llama 3 Community License Agreement.
+Appendix F Use of Large Language Models
+We employed ChatGPT to enhance the clarity of the manuscript by focusing on grammar corrections, shortening overly complex sentences, and providing alternative wording suggestions.
+All outputs were manually reviewed before inclusion, and no new technical material, code, results, or figures were generated by the tool.